### v2

# No Vulnerability Data, No Problem: Towards Predicting Mean Time To Remediate In Open Source Software Dependencies

Abstract.
Timely remediation of vulnerabilities in software dependencies is critical for the security of the software supply chain. As such, researchers have proposed tools and metrics to help practitioners assess the security practices of each of their dependencies. Conceptually, a dependency-focused Mean-Time-To-Remediate (MTTR) metric can provide a historical perspective on how long it takes a given package to update vulnerable versions of its dependencies. However, existing MTTR metrics focus on a package fixing bugs in its own code, not its dependencies. Simultaneously, existing dependency update metrics do not aggregate values for the entire package and are not sensitive to aspects important for vulnerabilities (e.g., floating version constraints). The goal of this study is to aid industry practitioners, including developers, in assessing the risk of dependencies through a novel metric approximating mean-time-to-remediate vulnerabilities in their dependencies that is evaluated by an empirical study. We propose a novel algorithm for computing MTTR called MTTRdep and a companion metric called Mean-Time-To-Updatedep (MTTUdep), which considers all version updates, including vulnerability fix updates. We conduct a large-scale study using $163,207$ packages in npm, PyPI, and Cargo, of which only $22,513$ packages produce MTTRdep because of the lack of vulnerability data. We further study how package characteristics (e.g., contributors and version counts) influence MTTUdep and MTTRdep and explore how long packages retain outdated vulnerable dependencies in npm, PyPI, and Cargo. Our results indicate that industry practitioners can reliably use MTTUdep as a proxy for MTTRdep when available vulnerability data is insufficient.
1. Introduction
Vulnerable dependencies are widely present in both open-source software (OSS) and proprietary codebases. According to the Synopsys 2025 “Open Source Security and Risk Analysis Report” (blackduck, ), 86% of codebases contain at least one vulnerable open source dependency, and 81% of codebases contain high or critical risk vulnerabilities resulting from dependencies. The majority of codebases had vulnerable dependencies for more than two years despite the availability of a fixed version (ossf-scorecard, ; blackduck-2021, ). This delay occurs because of fear of breaking changes and the cost associated with updating vulnerable dependencies to fixed versions (derrKeepMeUpdated2017, ).
Industry practitioners desire metrics that help assess the trustworthiness, risk, and security practices of open-source dependencies. For example, OpenSSF Scorecard (ossf-scorecard, ) considers 18 security practices to assess a package, including the “maintained” check, which determines if the project is maintained by checking activity in the last 90 days, and the “vulnerabilities” check detects if there are unfixed known vulnerabilities in the project or its dependencies.
Mean-Time-To-Remediate (MTTR) is also often discussed in the context of vulnerability management (sonatype-2019, ). In contrast to Scorecard’s ‘maintained’ metrics, MTTR provides a historical perspective. However, MTTR has largely been discussed in the context of a project fixing its own vulnerabilities, not vulnerabilities in dependencies (gokhaleTimeStructureBased1999, ; calvoApplyingGoalQuestion2023, ; johnsonSurveySoftwareTools1988, ; morrisonMappingFieldSoftware2018, ). Dependency update metrics such as technical lag (gonzalez-barahona_technical_2017, ) consider management of dependency versions; however, they have not been applied to vulnerabilities. Conceptually, a lower value in a dependency update metric suggests faster and more consistent updates, signaling ongoing maintenance and reliability (sonatype-2024, ), which could help assess the security risks of adopting a package. However, the connection between dependency updates and exposure to vulnerabilities in dependencies has not yet been demonstrated.
Practitioners assessing the risk associated with using a given dependency face two key challenges when considering how it manages vulnerabilities in its dependencies.
1. Insufficient Vulnerability Data. Currently, only 2.4K of npm packages have known vulnerabilities, while the ecosystem hosts over 3.27M packages. Practitioners can only compute MTTR for the dependents of those 2.4k packages.
2. Metric Limitations. The calculations used by existing dependency update metrics often do not handle floating version constraints or are not designed to handle vulnerabilities (coxMeasuringDependencyFreshness2015, ; gonzalez-barahona_technical_2017, ). Existing metrics also often calculate an update time for each of a dependency and for each version of a package, making it difficult to understand that package’s update practices as a whole. Finally, there are no existing publicly available dependency update metrics specifically for vulnerabilities. While some industry reports use MTTR in this way, their specific calculations are not available.
The goal of this study is to aid industry practitioners, including developers, in assessing the risk of dependencies through a novel metric approximating mean-time-to-remediate vulnerabilities in their dependencies that is evaluated by an empirical study. Specifically, we hypothesize that a package’s dependency update metric (Mean-Time-To-Update, MTTU) can serve as an indicator of how quickly the package would remediate security vulnerabilities (MTTR) in dependencies, even if the package never had a vulnerable dependency before. We also overcome challenges with dependency update calculations and propose novel algorithms for computing MTTU and MTTR, which we denote MTTUdep and MTTRdep, respectively.
We use MTTUdep and MTTRdep to answer the following research questions: RQ1:Can MTTUdep be used as a proxy for MTTRdep? RQ2:How do package characteristics influence MTTUdep and MTTRdep? RQ3:How long do packages have outdated dependencies, and how long do packages have vulnerable dependencies?
To answer these questions, we first collect package version release information and dependency relations of packages from the npm, PyPI, and Cargo ecosystems and compute MTTUdep and MTTRdep for $163,207$ packages. For RQ1, we perform an equivalence analysis using TOST (two one-side t-tests) and analyze correlations between MTTUdep and MTTRdep. For RQ1, we use correlation tests to understand the association between various package characteristics (e.g., contributors count, version count). We further conduct regression analysis to quantify which package characteristics matter more in influencing MTTUdep and MTTRdep values. Finally, for RQ1, we analyze the distributions of MTTUdep and MTTRdep to explore how long packages retain outdated and vulnerable dependencies in the ecosystems.
Contributions. In summary, this paper contributes (1) a detailed algorithm and process for quantifying the dependency update practice of a package using our novel metrics; (2) statistical hypothesis testing on using MTTUdep as a proxy for MTTRdep, when vulnerability data is not available; (3) a large-scale analysis of the dependency update metrics in npm, PyPI, and Cargo packages; (4) correlation and regression analysis illustrating which package characteristics impact the likelihood of higher MTTUdep and MTTRdep; and (5) actionable insights for open-source developers and supply chain security researchers based on our findings.
We provide our replication package in Zenodo (zenodo-artifact, ) (available for reviewers till Oct’25). Upon acceptance of the paper, we will make it public.
2. Related Work
In this section, we provide a brief overview of the related work.
Technical Lag. The concept of technical lag was first introduced by Gonzalez-Barahona et al. (gonzalez-barahonaTechnicalLagSoftware2017, ) for OSS packages. Essentially, “technical lag” quantifies how quickly software systems fall behind as new versions and updates are released. Zerouali et al. (zeroualiEmpiricalAnalysisTechnical2018, ) applied “technical lag” to the context of dependencies and conducted an empirical analysis of package dependency updates in the npm ecosystem. They found that outdated dependencies induce a median technical lag of 3.5 months in npm. Building on this, Decan et al.  (decanEvolutionTechnicalLag2018, ) conducted a longitudinal empirical study of ‘technical lag’ in the npm dependency network and explored how technical lag increases over time. They observed that technical lag for most npm packages increases during their lifespan, and technical lag occurred mainly due to the minor and patch releases of a dependency.
Further research by Zerouali et al. (zeroualiFormalFrameworkMeasuring2019, ) propose a formal framework for measuring technical lag in software ecosystems. They analyzed 4M releases of 500K npm packages considering the evolution of technical lag over time. They found that technical lag induced by direct dependencies in npm packages increases over time due to missed updates, including major releases. Stringer et al. (stringerTechnicalLagDependencies2020, ) study the technical lag of dependencies in a large-scale cross-ecosystem fashion containing packages from 14 package managers. They found that pinned dependencies are the main reason behind technical lag. Zerouali et al. (zeroualiMultidimensionalAnalysisTechnical2021, ) expand the idea of ‘technical lag’ into multiple dimensions (package lag, time lag, version lag, vulnerability lag, and bug lag) and study the technical lag in 140K Docker images. They found that the median time lag of community Docker images is over a year. Although previous studies have explored technical lag in terms of general dependency updates, understanding vulnerable dependency update practice was not their goal. In contrast, we study the dependency update metric and vulnerable dependency update metric and investigate their relationship.
Outdated and vulnerable dependencies. Kula et al. (kulaTrustingLibraryStudy2015, ) analyze the latency in adopting the latest version of a dependency in the Maven ecosystem. Their study found that developers are more likely to adopt the latest version for newly added dependencies than existing ones. In a follow-up study, Kula et al. (kula_developers_2018, ) examined library migration across GitHub projects and found that 81.5% of the projects keep using outdated dependencies. Cox et al. (coxMeasuringDependencyFreshness2015, ) introduce the concept of ‘dependency freshness’ to study dependency updatedness in the Maven ecosystem. They found that only $16.7\%$ of the dependencies display no update lag. Derr et al. (derrKeepMeUpdated2017, ) identify the root causes of outdated dependencies in the Android ecosystem and find that developers do not update their third-party library dependencies due to fear of breaking changes, lack of knowledge, and lack of motivation. Wang et al. (wangEmpiricalStudyUsages2020, ) conduct an empirical study on dependency update analysis on OSS packages and find that 50% of the packages use outdated dependencies. Huang et al. (huangCharacterizingUsagesUpdates2022, ) extend Wang et al. (wangEmpiricalStudyUsages2020, )’s study and find that one-third of the projects have a lag of one major version from the latest library version.
Pashchenko et al. (pashchenkoVulnerableOpenSource2018, ) studied the most used Java dependencies in SAP software and found that only updating the dependencies’ version can remove 81% vulnerable dependencies. Kula et al. (kulaDevelopersUpdateTheir2018, ) studied the update behavior of developers w.r.t. security advisories and found that developers do not update their vulnerable dependencies regularly. Kumar et al. (kumar_comprehensive_2024, ) conducted a study to understand how widespread vulnerabilities are and how quickly they are being fixed. They found that for most programming languages, a critical vulnerability persists on average for over a year before being fixed. Studies in outdated dependencies and vulnerable dependencies are focused on either all updates or only security updates but not both. In contrast, we focus on both outdated and vulnerable dependencies and explore their relationship using our proposed metrics.
MTTU and MTTR. The metrics MTTU and MTTR have been used in the software reliability and maintenance domain for a long time (gokhaleTimeStructureBased1999, ; calvoApplyingGoalQuestion2023, ; johnsonSurveySoftwareTools1988, ; morrisonMappingFieldSoftware2018, ). Researchers have also studied different security metrics (time to close bug/vulnerability, window of exposure, vulnerability count) (davidsonGoodBadUgly2009, ; beresUsingSecurityMetrics2009, ) of various categories (time metric, vulnerability metric) in the software security domain, but these metrics are focused on packages having the vulnerability. In our study, we focus on measuring dependency update metrics for packages having vulnerable dependencies, not the vulnerable package itself.
MTTR has also been used in different contexts in the industry, e.g., measuring the package’s security (mttr-dark-reading, ; mttr-plextrac, ) and measuring the package’s security in terms of dependency (sonatype-2019, ). The procedures for measuring MTTR and MTTU are often proprietary and not disclosed for academic research. For example, Sonatype’s 2024 report (sonatype-2024, ) measured MTTU and MTTR for Maven, but we could not reproduce it since the methodology is not available.
3. Challenges Applying Existing Update Metrics
In this section, we first provide an example of how the most prominent update metric with a published algorithm available in the literature, technical lag, works. Then, we describe the design gaps present in technical lag when measuring the updatedness of dependencies and why new metrics are needed.
Table 1. tLag measurement of codemod-cli package with one of its dependency, simple-git.
codemod-cli version: date simple-git constraint lastAllowed(simple-git) version: date latest(simple-git) version: date tLag
$0.8.6:\operatorname{2022-03-03}$ ${}^{\wedge}1.130.0$ $1.132.0:\operatorname{2020-03-12}$ $3.2.6:\operatorname{2022-02-17}$ $707$
$0.8.7:\operatorname{2022-03-05}$ ${}^{\wedge}2.48.0$ $2.48.0:\operatorname{2021-12-01}$ $3.2.6:\operatorname{2022-02-17}$ $78$
$0.9.0:\operatorname{2022-03-08}$ ${}^{\wedge}2.48.0$ $2.48.0:\operatorname{2021-12-01}$ $3.2.6:\operatorname{2022-02-17}$ $78$
$0.9.1:\operatorname{2022-03-17}$ ${}^{\wedge}2.48.0$ $2.48.0:\operatorname{2021-12-01}$ $3.3.0:\operatorname{2022-03-11}$ $100$
$0.9.2:\operatorname{2022-03-22}$ ${}^{\wedge}3.4.0$ $3.4.0:\operatorname{2022-03-18}$ $3.4.0:\operatorname{2022-03-18}$ 00
3.1. Technical Lag Using tLag
Time lag ($tLag$(zerouali_empirical_2018, ; decan_evolution_2018, ) is a way to measure technical lag to assess the outdatedness of a dependency in terms of time. Conceptually, $tLag$ measures the time difference between the release time of the latest version of a dependency and the release time of the version of the dependency used by a package, at the time the package was released.
Let $pkg$ be a package, and $pkg_{v}$ indicate specific version $v$ of $pkg$. Let $dep$ be a direct dependency of $pkg$. When specifying $dep$ as a dependency, $pkg_{v}$ specifies a dependency constraint $pkg_{v}.dep.\bar{c}$. Package managers use dependency constraints to resolve the highest version that satisfies the constraint. For example, a dependency constraint $pkg_{v}.dep.\bar{c}$ of “${}^{\wedge}1.2.3$” is satisfied for the range [¿=1.2.3, ¡ 2.0.0]. We use the function $lastAllowed(dep,\bar{c},t)$ to denote the version resolution of $dep$ for constraint $\bar{c}$ at time $t$. We further use the function $latest(dep,t)$ to denote the latest version of $dep$ at time $t$ and $time(\cdot)$ to denote the time a given version is released. Hence, the $tLag$ of $pkg_{v}$ for $dep$ is formally defined as follows:
(1) $tLag(pkg_{v},dep)=time(latest(dep,time(pkg_{v})))\\ -time(lastAllowed(dep,pkg_{v}.dep.\bar{c},time(pkg_{v})))$
Figure 1. Illustration of tLag, and MTTUdep/MTTRdep calculation using codemod-cli’s dependency relationship with simple-git. Blue time quantum indicates outdated dependency, red time quantum indicates outdated and vulnerable dependency, and yellow quantum indicates updated dependency.
Table 1 shows an example computation of technical lag for codemod-cli’s dependency on simple-git. The first two columns of Table 1 indicate the version with the date of different releases of package codemod-cli and the version constraint specified for dependency simple-git. The “lastAllowed(simple-git)” column, represents the resolved version of dependency simple-git with specified constraint at the time of the release of codemod-cli. For computing the technical lag, we need to know the latest available version of the dependency simple-git at the time of the release of each version of codemod-cli according to Eq. 1, which represents in column “latest(simple-git)”. The tLag column is the computed technical lag. We subtract the release date of lastAllowed(simple-git) from the release date of latest(simple-git) to obtain the technical lag for each release of codemod-cli. The computed tLag of five versions of codemod-cli were 707, 78, 78, 100, and 0, consecutively, in Table 1 and Fig. 1.
3.2. Design Gaps
In this section, we analyze the gaps available in tLag, which are used as motivators of the design choice of our proposed metrics.
1
Handling Floating Version Constraints. Floating version constraints are available in all major OSS ecosystems and are considered a good practice (he_pinning_2025, ; jafari_dependency_2023, ), as a package gets automatic updates whenever a newer version of the dependency is released. However, tLag does not sufficiently account for automatic updates that are allowed by floating version constraints. For example, tLag only checks the latest available version of the dependency against the package’s used dependency version when the package releases a new version. Even if the package releases a new version, and after that, the dependency releases a newer version that can be auto-updated by the package constraint, tLag cannot model this case in its design. Accounting for floating labels at the time of a vulnerability fix is particularly important when considering exposure to vulnerabilities, as an auto-update may automatically remediate a vulnerability.
2
Package-Level Metric. A dependency update metric should provide a way to aggregate all data points into one data for each package so that developers can make sense of the metrics and use them for comparison if desired. One of the findings of Wermke et al. (wermkeAlwaysContributeBack2023, ) was that developers’ dependency selection metrics and criteria focus on quickly accessible numbers and facts, such as downloads, GitHub stars, and time since last release, substantiating the necessity of having easily comprehensible metrics. However, tLag computes the technical lag for each version of the package for a dependency as shown in Table 1. For developers, understanding what to do with all of these tLag values for a package with multiple releases and multiple dependencies is hard. Having an aggregation would improve the understandability of dependency update metrics.
3
Handling Vulnerabilities. A security-oriented dependency update metric should incorporate the update characteristics of vulnerable dependencies. Technically, tLag’s measurement can be modified to only consider the dependency’s security fixes adopted by the package in its formal framework. However, no previous work measured tLag for measuring vulnerable dependency update practices.
4. Proposed Dependency Update Metrics
In this section, we first describe our design choices that fill the gaps presented in Section 3.2 for dependency update metrics measurement. Then, we provide an example of the nominal case using our metrics. We then provide a detailed methodology for defining the dependency update metrics in Section 4.3.
4.1. High-Level Overview
We design our metrics to only consider direct dependencies since the package only has control over which version to use for the direct dependencies. We begin by examining each of the $<$package, dependency$>$ relationships from a temporal perspective. We then split each $<$package, dependency$>$ relations into multiple intervals based on when the package or the dependency has a newer release (major, minor, or patch release). Our decision to split into intervals is to capture the benefits of floating version constraints without incentivizing this practice (Gap 1). [^1] To the best of our knowledge, no academic work quantifies optimal dependency specification (e.g., pinning vs floating version constraints) for balancing security benefits and the cost of maintenance. Therefore, our design does not discriminate between ways to specify dependency constraints (he_pinning_2025, ). Then, at each interval, the dependency constraint set by the package for the dependency is resolved with only the versions available at the beginning of the interval. We do this to make sure that the dependency resolution accounts for the historical version releases of the dependency.
After splitting into intervals, we mark “updated” as ‘false’ if the resolved dependency version for that interval does not match the highest available version of the dependency at that time and ‘true’ otherwise. Similarly, we mark “remediated” as ‘false’ if the resolved dependency version is vulnerable with a fixed version being available at the beginning of that interval and ‘true’ otherwise (Gap 3). We then aggregate the “updated” and “remediated” information of each $<$package, dependency$>$ to compute the MTTUdep and MTTRdep of each package (Gap 2).
4.2. Nominal Case With Our Metrics
Table 2. Running example of codemod-cli package with one of it’s depndency simple-git.
row pkg pkg version dep dep constraint dep version dep highest rel. Interval start Interval end updated remediated
1 codemod-cli $0.4.0$ simple-git ${}^{\wedge}1.130.0$ $1.132.0$ $2.13.1$ $\operatorname{2020-07-16}$ $\operatorname{2020-07-17}$ false true
$\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$
82 codemod-cli $0.8.6$ simple-git ${}^{\wedge}1.130.0$ $1.132.0$ $3.2.6$ $\operatorname{2022-03-03}$ $\operatorname{2022-03-05}$ false true
83 codemod-cli $0.8.7$ simple-git ${}^{\wedge}2.48.0$ $2.48.0$ $3.2.6$ $\operatorname{2022-03-05}$ $\operatorname{2022-03-08}$ false true
84 codemod-cli $0.9.0$ simple-git ${}^{\wedge}2.48.0$ $2.48.0$ $3.2.6$ $\operatorname{2022-03-08}$ $\operatorname{2022-03-11}$ false true
85 codemod-cli $0.9.0$ simple-git ${}^{\wedge}2.48.0$ $2.48.0$ $3.3.0$ $\operatorname{2022-03-11}$ $\operatorname{2022-03-17}$ false false
86 codemod-cli $0.9.1$ simple-git ${}^{\wedge}2.48.0$ $2.48.0$ $3.3.0$ $\operatorname{2022-03-17}$ $\operatorname{2022-03-18}$ false false
87 codemod-cli $0.9.1$ simple-git ${}^{\wedge}2.48.0$ $2.48.0$ $3.4.0$ $\operatorname{2022-03-18}$ $\operatorname{2022-03-22}$ false false
88 codemod-cli $0.9.2$ simple-git ${}^{\wedge}3.4.0$ $3.4.0$ $3.4.0$ $\operatorname{2022-03-22}$ $\operatorname{2022-03-29}$ true true
$\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$
In this section, we explain how to calculate the dependency update metrics. As a running example, we consider the dependency relation between codemod-cli and simple-git, as shown in Table 2.
We split the dependency relations into multiple intervals based on the release of a new version of the package “codemod-cli” or dependency “simple-git”. In each interval, for resolving the dependency constraint by the package codemod-cli, we only consider the dependency versions available at the beginning of the interval (“interval start”). From $V0.4.0$ to $V0.8.6$ (row 1-82), codemod-cli has the dependency constraint simple-git ${}^{\wedge}1.130.0$ which resolves into $V1.132.0$. simple-git has the highest release $V2.13.1$ at $\operatorname{2020-07-16}$ (row 1) and $V3.2.6$ at $\operatorname{2022-03-03}$. As a result, codemod-cli does not have the highest available release of simple-git in these intervals (rows 1-82). In $V0.8.7$, codemod-cli updated the constraint for simple-git to ${}^{\wedge}2.48.0$. Even with this update, codemod-cli has not changed the constraint of simple-git to the available highest major version 3 and rather stayed at major version 2. codemod-cli have ${}^{\wedge}2.48.0$ constraint for simple-git from $V0.8.7$ to $V0.9.1$. The false value in the “updated” column means that the package has an outdated version of the dependency in this interval.
At the beginning of row 85, the resolved version of simple-git $V2.48.0$ was found vulnerable to four vulnerabilities (CVE-2022-24066 (CVE-2022-24066, ), CVE-2022-24433 (CVE-2022-24433, ), CVE-2022-25912 (CVE-2022-25912, ), and CVE-2022-25860 (CVE-2022-25860, )). In this example, we only consider CVE-2022-24433 as a vulnerability, which was fixed in $V3.3.0$, released on $\operatorname{2022-03-11}$. For this reason, we mark the intervals from rows 85-87 as “remediated” = false since codemod-cli has a vulnerable version of simple-git even though a fixed version is available. The false value in the “remediated” column means that the package has an outdated and vulnerable version of the dependency in this interval. In $V0.9.2$, codemod-cli bumped the constraint of simple-git to ${}^{\wedge}3.4.0$ which resulted in $V3.4.0$, a fixed version of the above vulnerability. So codemod-cli has the highest available version of simple-git in this interval (row 88) and so marked as true in the “remediated” column.
When computing the Time-To-Update (TTUdep) for dependency relation between codemod-cli and simple-git, we sum up the intervals marked false in the “updated” column. For codemod-cli, the TTUdep is the delta between $\operatorname{2020-07-16}$ to $\operatorname{2022-03-22}$ (rows 1 - 87) which is $614$ days. Similarly, for computing Time-To-Remediate (TTRdep), we sum up the intervals marked false in the “remediated” column. For codemod-cli, the TTRdep is the delta between $\operatorname{2022-03-11}$ to $\operatorname{2022-03-22}$ (rows 85 - 87) which is $11$ days.
4.3. Metrics Definitions
We present a formal definition of our proposed metrics in this section.
Metric 1 ($\operatorname{\mathbf{Mean-Time-To-Update}}_{dep}$ : $\mathbf{MTTU}_{dep}$).
MTTUdep of a package is the average aggregated time the package uses an outdated direct dependency version in its lifetime.
We have described how to calculate the TTUdep for a $<$package, dependency$>$ relationship in Section 4.2. Formally speaking, TTUdep of a package $p_{i}$ with considering only dependency $p_{j}$ ($<p_{i},p_{j}>$ relationship) is:
(2) ${TTU}_{dep}(p_{i},p_{j})=\sum\operatorname{intervals\ with\ updated=False}$
MTTUdep of package $p_{i}$ with $n$ direct dependencies is:
(3) ${MTTU}_{dep}(p_{i})=\frac{\sum\limits_{j=1}^{n}{TTU}_{dep}(p_{i},p_{j})}{n}$
Metric 2 ($\operatorname{\mathbf{Mean-Time-To-Remediate}}_{dep}$ : $\mathbf{MTTR}_{dep}$).
MTTRdep of a package is the average aggregated time a package uses an outdated and vulnerable direct dependency version in its lifetime.
We have described how to calculate the TTRdep for a $<$package, dependency$>$ relationship in Section 4.2. Formally speaking, TTRdep of a package $p_{i}$ with considering only dependency $p_{j}$ ($<p_{i},p_{j}>$ relationship) is:
(4) ${TTR}_{dep}(p_{i},p_{j})=\sum\operatorname{intervals\ with\ remediated=False}$
MTTRdep of package $p_{i}$ with $n$ direct historically vulnerable dependencies is:
(5) ${MTTR}_{dep}(p_{i})=\frac{\sum\limits_{j=1}^{n}{TTR}_{dep}(p_{i},p_{j})}{n}$
5. Methodology
In this section, we first present vulnerability, package metadata (versions and dependency relations), and package characteristics collection process. We then present a statistical testing process to verify if MTTUdep can be a proxy for MTTRdep (RQ1). We then present correlation tests and regression analysis to understand how package characteristics impact MTTUdep and MTTRdep (RQ1).
5.1. Data Collection
5.1.1. Vulnerability Information
We collect the CVE data for our chosen three ecosystems from osv.dev (osv-dev, ) for npm, PyPI, and Cargo packages on 2024-09-12. We rely on OSV since OSV aggregates CVE data from multiple sources (e.g., GitHub Security Advisories, PyPA, GoVulDB) (osv-data-sources, ) in one place. After downloading JSON formatted CVE data from OSV, we convert it into an SQL table that includes ecosystem, package name, CVE ID, version where the vulnerability was introduced, and version where the vulnerability is fixed.
5.1.2. Package Metadata
We collect the package-version data and dependency information for npm, PyPI, and Cargo packages from deps.dev on 2024-08-20, similar to other previous studies (lin_untrustide_2024, ; zahan_software_2023, ). In our dataset, we have initially $2,603,314$ npm, $274,720$ PyPI, and $122,069$ Cargo packages. We collected data from deps.dev since it provides all the package versions and dependency information for our three chosen ecosystems. After data collection, we split each $<$package, dependency$>$ relation into multiple intervals. Since the dependency resolution can be complex and differs across ecosystems, we use deps.dev to perform the dependency resolution allowed by the dependency constraints with our added requirement of only using the available versions of the dependency before the interval start time.
5.1.3. Package Characteristics
We examined several characteristics (from Saini et al. (saini_investigating_2020, )) for packages: the number of contributors; the number of dependencies and dependents; the number of version releases; the SourceRank score, and the number of forks and stars. On 2025-01-11, we downloaded these characteristics of each package from libraries.io to understand if these factors correlate with package’s MTTUdep and MTTRdep values. Libraries.io is used by other research as a data source (gu_self-admitted_2023, ; cao_towards_2023, ; he_migrationadvisor_2021, ). When counting the number of dependencies, we only use the number of dependencies of packages in their latest version. To ensure construct validity in downloading these factors, we manually inspected a sample of packages to verify the factors, and we found the downloaded factors to be accurate. Additionally, we computed the number of major versions and package ages for our packages and added them to our analysis.
5.2. Package Inclusion Exclusion Criteria
We begin with an initial dataset of $3,000,103$ ($2,603,314$ npm, $274,720$ PyPI, and $122,069$ Cargo) packages collected from deps.dev. Our first step is to apply two inclusion criteria: (1) the package must be at least two years old (operationalized by the difference between the first and last version release); and (2) the package must have at least one residual activity (e.g., one version release) in the last two years. Miller et al. (miller_understanding_2025, ) used two years of residual activity followed by two years of no maintenance as criteria to find out the abandoned packages. Our criteria are inspired by Miller et al. since we want to include the packages that are maintained. Moreover, “two years” is a commonly used standard adopted by other research to measure whether a package is maintained or not (li_comparison_2023, ). However, our package selection criteria might miss packages that are less than two years old or have had no activity in the last two years (e.g., feature complete packages (coelho2017modern, )) even if they are not abandoned. We then used our exclusion criteria: a package without any dependency should be excluded. Since our metrics characterize packages’ dependency update practice, packages without any dependencies do not fit into our study. The number of packages after this exclusion criteria is $163,207$ ($117,129$ npm, $42,777$ PyPI, and $3,301$ Cargo packages). Out of these $163,207$ packages, $22,513$ ($17,263$ npm, $5,158$ PyPI, and $92$ Cargo) packages have at least one vulnerable dependency in their lifetime, supporting our initial motivation of lack of vulnerability data.
5.3. Metrics Implementation
After resolving the dependencies and applying our inclusion-exclusion criteria, we store the data in a PostgreSQL database. Given the large size of the dataset, working directly with the raw data would be impractical. We also create indexes on the most frequently accessed keys to speed up data retrieval for our metrics calculations.
In the database, we compute if the resolved dependency version matches the highest available version of the dependency during each interval and mark each interval accordingly, as shown in the “updated” column in Table. 2. Similarly, we then compute if the resolved dependency version for each interval is vulnerable to any security advisory even though a fixed version is available and mark the interval accordingly in the “remediated” column in Table. 2.
5.4. Statistical and Correlation Tests
To answer RQ1, we claim that if MTTUdep and MTTRdep are statistically equivalent, then MTTUdep can serve as a proxy for MTTRdep. Since we did not expect our samples to follow a normal distribution, we initially selected the non-parametric Mann-Whitney test. However, we also applied the t-test to ensure a more conservative analysis, as our data is not ordinal. Both tests were conducted using the Python library scipy, and the Wilcoxon test was also implemented using the same package. To assess normality, we used the Kolmogorov-Smirnov test.
Our initial hypotheses focused on the equivalence of data distributions. For these, following the approach of Papotti et al. (papotti_acceptance_2024, ), we employed the TOST (Two One-Sided Test) approach, originally proposed by Schuirmann et al. (schuirmann_hypothesis-testing_1981, ) and commonly used in pharmacological and food sciences to evaluate equivalence within a specified range (guidance_statistical_2001, ; meyners_equivalence_2012, ). Other research also used the TOST test in their analysis (shelton_rosita_2021, ; labunets2018no, ; labunets2017graphical, ; palheiros2024does, ). According to TOST, two distributions, $x$ and $y$, are considered equivalent if $x\cdot\delta<y<x\cdot 1/\delta$, where $\delta=0.8$. The TOST method can use either Mann-Whitney test or t-test as its underlying direction test. We reported the results of the t-test because it was more conservative than the Mann-Whitney test.
Additionally, to measure the association between MTTUdep and MTTRdep, we calculated both the Pearson correlation coefficient and Spearman’s rank correlation coefficient (similar to Cox et al. (cox_measuring_2015, )). High ($>0.7$) or moderate coefficient values ($>0.5$) would suggest a strong to moderate positive relationship (ratner2009correlation, ). Such findings would also support the notion that MTTUdep could effectively serve as a proxy for MTTRdep.
5.5. Package Characteristics Analysis
For analyzing eight package-level characteristics in RQ1, we use correlation tests to estimate if there is any correlation between these characteristics and packages’ corresponding MTTUdep and MTTRdep values. Specifically, we test a hypothesis about the association between the following independent variables and their continuous dependent variables (MTTUdep and MTTRdep):
Contributors Count ($H_{1}$). We hypothesize packages with fewer contributors are less likely to have updated dependencies since they have less capacity for maintenance and dependency management.
Dependents Count ($H_{2}$). We hypothesize packages with fewer dependents are less likely to have updated dependencies. Fewer dependents may lead to fewer feature updates, fewer bug fixes, and fewer version releases, which in turn may result in less updated dependencies.
Dependency Count ($H_{3}$). We hypothesize packages with fewer dependencies are more likely to have updated dependencies. We expect fewer dependencies to be more manageable, and thus, these projects may have more updated dependencies.
Version Count ($H_{4}$). We hypothesize packages with fewer number of version releases are less likely to have updated dependencies. Since these packages have less available versions, they may pay less attention to their dependency management.
Major Version Count ($H_{5}$). We hypothesize packages with fewer major version releases are more likely to have updated dependencies. Since these packages have fewer major versions to maintain, they may pay more attention to their dependency management.
Package Age ($H_{6}$). We hypothesize packages with lower ages are more likely to have updated dependencies. Lower package age may mean the development team is more proactive in dependency management since the package is not old or mature enough. Lower package age may also mean that the dependencies do not have scope for publishing many newer versions, which in turn may mean more updated dependencies in these packages.
SourceRank ($H_{7}$). The Package SourceRank score indicates the package quality, popularity, and community metrics calculated in libraries.io dataset (saini_investigating_2020, ; sun_using_2023, ). This metric depends on several factors, such as the presence of a README file, license, following SEMVER, recent updates, and the number of contributors. We hypothesize packages with lower SourceRank scores are less likely to have updated dependencies. Since these packages are of low quality, they may not have a lot of dependents or contributors, which in turn may result in less updated dependencies.
Forks And Stars Count ($H_{8}$). We hypothesize packages with fewer forks and stars are less likely to have updated dependencies. Fewer forks and stars may mean fewer people are using and looking into these packages, which in turn may mean less activity, less version releases, and less updated dependencies in these packages.
5.6. Regression Analysis
Correlation results indicate whether two independent variables are associated with each other or not. However, correlation does not establish causation. To identify such a cause-and-effect relationship using mathematical modeling, we opt for applying a technique of multilinear regression analysis. The multilinear regression model takes independent variables (e.g., package characteristics) and one dependent variable (MTTUdep or MTTRdep) and gives results on the relationship between each independent variable and the dependent variable. This analysis provides valuable insights on predicting MTTUdep or MTTRdep using the package’s characteristics. This analysis also results in $p$-values and coefficients, indicating which package characteristics might have a significant impact on MTTUdep or MTTRdep.
This process will produce a $p$-value for each independent variable to indicate whether the relationship between this variable and the outcome is statistically significant. Finally, we consider the variables with $p$-value $<0.05$ as significant variables, and the package characteristics associated with these variables are identified as key package characteristics. We use statsmodels package of Python to conduct the correlation tests and to build the regression model.
6. Results
6.1. RQ1: Can MTTUdep serve as a proxy for MTTRdep?
We statistically assess the relationship between MTTUdep and MTTRdep to determine if MTTUdep can effectively serve as a proxy for MTTRdep in  RQ1.
We used TOST (Two One-Sided Test) to investigate the equivalence of MTTUdep and MTTRdep. The single one-sided tests are done with the Mann-Whitney test. The results of the TOSTs are reported in Table 3. We also visualized the TOST for MTTUdep and MTTRdep from all three ecosystems in Fig. 2.
Table 3. TOSTs Results
Ecosystem MTTUdep ×$\times 0.8$ $<$ MTTRdep $<$ MTTUdep ÷$\div 0.8$
[…]
The green area corresponds to $x\cdot\delta<y<x\cdot 1/\delta$ where $\delta=0.8$, $x$ is MTTUdep, and $y$ is MTTRdep.
Figure 2. Distribution behind TOST MTTUdep vs. MTTRdep.
The results of the TOSTs show statistical equivalence between MTTUdep and MTTRdep ($p<0.05$) within the given range. Further, we also ran correlation analyses between MTTUdep and MTTRdep, in which both Spearman and Pearson correlation coefficients are $0.5$. The correlation analysis supports the results of our TOST tests as we can conclude that MTTUdep and MTTRdep are moderately and positively correlated since all correlation coefficients are $>=0.5$.
This result confirms that as the MTTUdep increases for a package, it is highly probable that the MTTRdep also increases (and vice versa), supporting the use of MTTUdep as a proxy for MTTRdep when we do not have enough vulnerability data to compute MTTRdep.
Takeaway 1: MTTUdep can be used as a proxy for MTTRdep since MTTUdep and MTTRdep are statistically equivalent (and moderately and positively correlated) in all three observed ecosystems.
6.2. RQ1: How do other package characteristics influence MTTUdep and MTTRdep?
Figure 3. Correlation matrix on packages’ characteristics and MTTUdep and MTTRdep values.
6.2.1. Correlation Analysis.
We tested eight package-level characteristics that might influence MTTUdep and MTTRdep. We present the correlation matrix with the packages’ characteristics and MTTUdep and MTTRdep data in Figure 3. The correlation heatmap provides insight into the relationships between various factors and their potential impact on MTTUdep and MTTRdep. The results indicate weak correlations between MTTUdep and the other factors.
Contributors Count ($H_{1}$): We hypothesized that packages with fewer contributors would be less likely to maintain updated dependencies. From the heatmap, we can see that the contributors count show minimal correlations with MTTUdep (-0.04) and MTTRdep (-0.07), suggesting a limited influence on dependency management.
Dependents Count ($H_{2}$): We hypothesized that packages with fewer dependents are less likely to have updated dependencies. The results indicate a negligible correlation between dependents count and MTTUdep or MTTRdep. This suggests that the number of dependents does not play a substantial role in determining the timeliness of dependency updates.
Dependency Count ($H_{3}$): We hypothesized packages with fewer dependencies to have more updated dependencies. The results show almost no correlation between dependency count and MTTUdep or MTTRdep. This implies that the number of dependencies in a package is not associated with dependency management efficiency.
Version Count ($H_{4}$): We hypothesized that packages with fewer version releases are less likely to have updated dependencies. However, the correlation matrix revealed no meaningful relationship between the number of version releases and either MTTUdep or MTTRdep. This suggests that the frequency of version releases does not significantly impact how quickly dependencies are updated.
Major Version Count ($H_{5}$): We hypothesized that packages with fewer major version releases are more likely to have updated dependencies. However, the correlation matrix revealed no meaningful relationship between the number of major version releases and either MTTUdep or MTTRdep. This suggests that the number of available major versions to maintain does not significantly impact how quickly dependencies are updated.
Package Age ($H_{6}$): We hypothesized that packages with lower ages are more likely to have updated dependencies. However, the correlation matrix invalidated this hypothesis with a weak correlation score ($0.32$). This suggests that as packages become mature with age, they tend to keep dependencies less updated. However, this conclusion should be treated with caution since our metrics might be age-sensitive (more discussion on this in Section 7).
SourceRank ($H_{7})$. We hypothesized that low quality packages (e.g., packages with lower SourceRank scores) are less likely to have updated dependencies. The correlation matrix supports this hypothesis for MTTRdep with a weak correlation score ($-0.22$), which means low-quality packages have higher MTTRdep values. However, low quality packages do not necessarily have higher MTTUdep values ($-0.06$ coefficient).
Forks And Stars Count ($H_{8}$): We hypothesized that packages with fewer forks and stars are less likely to have updated dependencies. The results reveal weak to moderate positive correlations between stars (0.65) and forks (0.72) with maintainers count, indicating that more popular packages may have more active development. However, their relationships with MTTUdep and MTTRdep remain insignificant, suggesting that forks and stars do not directly translate into faster dependency updates.
Overall, our results indicate weak or negligible correlations between the tested factors and dependency management efficiency as measured by MTTUdep and MTTRdep. These findings suggest that the analyzed factors may not strongly influence how updated dependencies are, and further exploration of other variables or non-linear relationships is recommended.
Takeaway 2: Package characteristics (except package age and SourceRank score) have negligible correlations with dependency updatedness. Older packages are associated with higher MTTUdep, and popular and better quality packages (higher SourceRank score) are associated with lower MTTRdep.
6.2.2. Regression Analysis (MTTUdep As Dependent Variable).
Interpreting the multilinear regression model’s characteristics, we found 0.073 as R-squared value when MTTUdep was the dependent variable. R-squared value 0.073 indicates that 7.3% variation in the dependent variable can be explained by the model. Although the model is statistically significant (p-value ¡ 0.05 from the F-test), lower R-squared value indicates that other factors, beyond the nine dependent variables, also substantially impact MTTUdep. The F-statistic of this model is large (1399), and Prob(F-statistic) is 0.00. This indicates that at least one of the independent variables (or predictors) has a non-zero relationship with MTTUdep.
We then look into the coefficients and p-values associated with each of the independent variables. The coefficient indicates the expected change in the dependent variable for a one-unit change in one independent variable while holding other independent variables constant. The p-value associated with the t-statistic indicates whether the independent variable is statistically significant in explaining the variation in the dependent variable. A lower p-value (¡ 0.05) would indicate statistical significance.
We have found that some independent variables have positive coefficients, which indicates an increase in one of these variables would result in an increase in our dependent variable, MTTUdep. Independent variables with positive coefficients are contributors count (0.0204), stars (0.0035), dependencies count (2.3027), and package age (0.1257). The dependencies count has the largest positive coefficient, which means that packages with more dependencies need longer times to update their dependencies. In addition, older packages tend to have longer MTTUdep as indicated by the 0.1257 coefficient, justifying our hypothesis ($H_{3}$) on dependencies count.
We also found some independent variables have negative coefficients. Negative coefficients indicate that an increase in that independent variable would result in a decrease in the dependent variable, MTTUdep. Independent variables with negative coefficients are dependents count (-0.0013), forks (-0.0085), SourceRank (-24.3203), and major versions count (-0.9068). SourceRank, which indicates package quality and popularity, shows strongly negative coefficients. Higher SourceRank indicates an overall better project. The model predicts if SourceRank for a package increases by 1 (moving to a “better” rank), MTTUdep drops by 24 days. Moreover, packages that have released more major versions tend to have a lower MTTUdep. Finally, all dependent variables except dependent count were found to be statistically significant (p-value ¡ 0.05).
Takeaway 3: Packages with a lot of dependencies or older packages are more likely to have higher MTTUdep. Also, popular and better quality packages (with higher SourceRank scores) are more likely to have lower MTTUdep.
6.2.3. Regression Analysis (MTTRdep As Dependent Variable).
For MTTRdep as the dependent variable, the multilinear regression model results in 0.097 as R-squared value. R-squared = 0.097 indicates that the model explains 9.7% of the variation in MTTRdep. A lower R-squared value indicates that other unmodeled factors play an important role in determining MTTRdep, which is similar to our observations in MTTUdep. The F-statistic of this model is 246.8, and Prob(F-statistic) is 0.00. This indicates the model is statistically significant and at least one of the dependent variables (or predictors) has a non-zero relationship with MTTRdep.
We then look into the coefficients and p-values associated with each of the independent variables. We found that forks, stars, SourceRank, dependencies count, major version count and package age are statistically significant (p-value ¡ 0.05). On the other hand, the contributors count, and dependents count are not statistically significant (p-value ¿ 0.05) in modeling MTTRdep.
As a predictor, forks (coefficient 0.0070) and stars (coefficient 0.0014) indicate a positive but small effect. More forks and stars are associated with slightly higher MTTRdep. Dependencies count (coefficient 0.5028), and package age (coefficient 0.1239) show a moderate positive effect on MTTRdep. Additionally, the major version count (coefficient -2.4954) shows a moderate negative effect, whereas SourceRank (coefficient -34.8183) shows a strong negative effect on MTTRdep. This observation is similar to our previous observation with MTTUdep. According to the model’s prediction, an increase in SourceRank by 1 (i.e., a higher SourceRank indicates a better package) would result in a 34.8-day reduction in MTTRdep.
Takeaway 4: Similar to MTTUdep, older packages are more likely to have higher MTTRdep. Also, popular and better quality packages (with higher SourceRank scores) are more likely to have lower MTTRdep. Packages with higher major versions are more likely to have lower MTTRdep.
6.3. RQ1: How long do packages have outdated dependencies, and how long do packages have vulnerable dependencies?
In RQ1, we empirically analyze the MTTUdep and MTTRdep metrics for the three ecosystems. We choose violin plot instead of box-plot since violin plot shows everything box-plot shows, e.g., medians, ranges, variability, and violin plot’s shape shows the density of the data similar to a density estimation plot (hintzeViolinPlotsBox1998, ).
Figure 4. ${MTTU}_{dep}$
Figure 5. ${MTTR}_{dep}$
Fig. 4 shows a violin plot of MTTUdep in npm, PyPI, and Cargo. All the plots are right-skewed, indicating that most packages have a low MTTUdep. For instance, $50\%$ npm package has MTTUdep of less than $174$ days. In terms of distribution width, Cargo has a compact distribution, and npm is the most widely spread (Cargo ¡ PyPI ¡ npm). Also, every plot has a long tail, indicating that every ecosystem has some packages that do not update their dependencies for a long time (max MTTUdep $=4296$ days).
Fig. 5 shows a violin plot visualizing the distribution across MTTRdep in days for Cargo, npm, and PyPI packages. The first, second (mean), and third quantiles are shown as , , and , respectively. The overall pattern is similar to MTTUdep. The relative distribution width and the right-skewed nature of MTTRdep is similar to MTTUdep. Cargo and PyPI have concentrated distributions, with most MTTRdep (e.g., interquartile range) being less than $219$ and $281$ days. Conversely, npm shows a more spread distribution than Cargo and PyPI, indicating variability in MTTRdep in npm.
We could compute MTTUdep for $163,207$ ($117,129$ npm, $42,777$ PyPI, and $3,301$ Cargo) packages, and MTTRdep for $22,513$ ($17,263$ npm, $5,158$ PyPI, and $92$ Cargo) packages. This corroborates our initial motivation for conducting this study, the lack of vulnerability data.
Takeaway 5: Even without potentially abandoned packages, a subset of packages show extreme delays in keeping their dependencies and vulnerable dependencies updated.
Comparison to Prior Work. Similar to prior research (kula_developers_2018, ), we also found that the majority of packages in an ecosystem do not keep dependencies updated. We also observed a long-tail distribution for both MTTUdep and MTTRdep in all three ecosystems. A significant number of packages lag behind in keeping dependencies updated, which is similar to the observations of Cox et al. (cox_measuring_2015, ). In short, regularly updating dependencies is not a widespread practice. Our observation is similar to vulnerable dependency updates as well. Many packages took months or longer to remediate known vulnerable dependencies, similar to observations from Kula et al. (kula_developers_2018, ). This suggests that challenges in keeping dependencies updated are not specific to any ecosystem but rather general in nature.
7. Discussions And Implications
When MTTUdep and MTTRdep should be used. Our study provides new insights by establishing the relationship between MTTUdep and MTTRdep. We found that although a package’s MTTUdep is moderately correlated with its MTTRdep, MTTUdep and MTTRdep are statistically equivalence. In practical terms, packages that are slow to update dependencies also tend to be slow in updating vulnerable dependencies. This finding implies that improving general dependency update practice will likely improve vulnerable dependency update practice, which seems logical, but we provide statistical evidence supporting our hypothesis. Having said that, developers should use MTTRdep when available and only MTTUdep as a proxy for MTTRdep when MTTRdep is not available.
Package Age’s Influence In MTTUdep and MTTRdep. Package age is weakly correlated with MTTUdep and has a negligible correlation with MTTRdep in our analysis. In our regression analysis, we have used package age as an independent variable with MTTUdep and MTTRdep as dependent variables. However, since we design our metrics based on a temporal perspective, our metrics could have an inherent age sensitivity. So, conclusions like ‘older packages tend to have slightly higher MTTUdep’ should be treated cautiously. Since older packages might have more data points (negligible but positive correlation with version count, major version count, and dependencies count), package age could influence our analysis. Future research should explore how to model the package age in MTTUdep and MTTRdep (e.g., one way could be to normalize the metrics by package age or by only taking the last $n$ years into account).
Influence of Unmodeled Package Characteristics. Prior work shows that developers hesitate to update dependencies due to a fear of breaking changes, a lack of awareness or knowledge about available updates, and sometimes a lack of motivation to invest time in updates (derrKeepMeUpdated2017, ). These human factors are the likely reasons behind the long tail of MTTUdep and MTTRdep data. Our findings also support this. Packages with many contributors or higher dependents count do not have a better MTTUdep or MTTRdep, in contrast to our hypothesis. A package used by thousands of other packages is not guaranteed to have dependencies updated. Both multilinear regression models in our study present an R-squared value of less than 10%. A lower R-squared value indicates that other unmodeled factors substantially affect MTTUdep and MTTRdep. The development team dynamic, the use of pinning and floating, the cost and efforts needed in testing, and the size of the codebase could be such possible unmodeled factors.
Practical implications for developers. We confirmed that improving general dependency update frequency likely improves security as well. This should be a strong incentive for the developers to treat dependency updates as an important part of routine maintenance, not an afterthought. Even scheduling periodic dependency update sprints or using automated update notifications to stay on top of the new releases can be an effective strategy. Allowing the latest version of the dependency using the dependency specification (e.g., * or $latest$) is the best way to keep MTTUdep and MTTRdep low (even zero). However, that might not be possible due to the issues with breaking changes (ochoa_breaking_2022, ; raemaekers_semantic_2014, ; venturini_i_2023, ). Allowing auto updates of minor and patch releases by the dependency could help reduce the MTTUdep and MTTRdep. This ensures developers get incremental improvements and fixes without manual effort. Similarly, pinning should be avoided since pinning does not allow any auto-updates, which will start increasing the MTTUdep as soon as the dependency releases a new version (even a patch release). Whenever a vulnerability is found in the dependency and a fixed version is available, developers should prioritize remediating that vulnerability. Finally, a mixed strategy with automatic minor and patch updates, alongside manual major version updates, could be the most effective strategy to keep the MTTUdep and MTTRdep minimal. Our recommendation relies on developers using SEMVER correctly (pinckney_large_2023, ; li_large-scale_2023, ; dietrich_dependency_2019, ) so that the benefits of floating versions can be leveraged.
Gaming Metrics. While it might seem that using a loose floating constraint and bumping dependency constraint periodically could “game” our metrics, the underlying outcome is that a package spends minimal time with outdated dependency versions and ensures that security fixes are adopted automatically. By doing so, this strategy reduces the risk of exploitation from using an outdated dependency version with known security vulnerability, such as log4Shell (everson_log4shell_2022, ; rahman_less_2024, ). In short, trying to game our metrics by adopting floating version constraints is essentially reducing this attack vector for a package. However, this approach also comes with a tradeoff. This strategy might make the packages susceptible to malicious package updates (xz-incident (williams_research_2025, )). Since floating version constraints allow automatic updates, it can also facilitate propagating malicious package updates.
Influence of Dependencies Count in MTTUdep and MTTRdep. In our design, we compute MTTUdep and MTTRdep by dividing the summation of TTUdep and TTRdep by the number of dependencies and number of historically vulnerable dependencies, respectively, in Eq. 3 and  5. A package with a high number of dependencies has more room for having a better value in MTTUdep while keeping the majority of dependencies updated. However, MTTRdep is not influenced by dependencies count since we only consider the amount of time spent with vulnerable dependency versions and the number of historically vulnerable dependencies.
8. Threats to Validity
In this section, we discuss potential limitations and how they might impact the interpretation of our results.
External Validity. The main external validity threat is the generalizability of our results to characterize other ecosystems. We carefully chose the three ecosystems to have a diverse set of ecosystems, where npm is the largest, PyPI is the oldest, and Cargo is the newest among the major software ecosystems. While each ecosystem possesses unique features that might not directly correlate with those we studied, we believe the insights gained should also be broadly applicable to other ecosystems.
Internal Validity. We use the security advisory dataset from OSV.dev, which may not be comprehensive. If an advisory is published but not included in the OSV dataset, that may impact our results. Additionally, we do not consider whether the vulnerable dependency version is exploitable or reachable (cisa-vex, ) from the package. We treat all vulnerabilities equally regardless of the CVSS score or the severity of the vulnerability. After downloading the data from deps.dev, we manually checked 20 packages’ versions and relations with the public package registries and found that the data is accurate. Moreover, each package manager has its own way of handling dependency resolutions, and for the dependency resolutions, we rely on the Open Source Insights (depsdev, ) resolved version data. Our analysis omits package versions not adhering to SEMVER rules, a conservative choice to enable a more rigorous analysis. In addition, Open Source Insights dependency resolution fails in some cases (e.g. missing timestamp), and we mark those as warnings in our dataset. We do not calculate update metrics for those packages, and we argue that this might have a very small impact on our results. In addition, we only consider runtime dependencies in our MTTUdep and MTTRdep analysis and omit dev and optional dependencies. Additionally, some dependencies might be more important than others depending on the context; however, we treat each dependency equally since modeling dependencies’ importance is out of the scope of our study.
9. Conclusion and Future Works
In this study, we introduced two dependency update metrics, MTTUdep and MTTRdep, to quantify the updatedness of dependencies in open-source software packages. Our large-scale empirical analysis across the npm, PyPI, and Cargo ecosystems demonstrated that MTTUdep can serve as a proxy for MTTRdep when vulnerability information is unavailable. Furthermore, our statistical analysis highlighted the relationships between package characteristics and dependency update behavior, providing actionable insights for developers, maintainers, and software supply chain researchers. Future research can explore additional factors influencing dependency update practices, such as the severity of vulnerabilities, organizational policies, or developer incentives. Expanding this analysis to other ecosystems, combining with transitive dependencies, and incorporating qualitative insights from developers could further refine our understanding of dependency updatedness.

### v3

@@
-No Vulnerability Data, No Problem: Towards Predicting Mean Time To Remediate In Open Source Software Dependencies
+How Quickly Do Development Teams Update Their Vulnerable Dependencies?
 Abstract.
-Timely remediation of vulnerabilities in software dependencies is critical for the security of the software supply chain.
-As such, researchers have proposed tools and metrics to help practitioners assess the security practices of each of their dependencies.
-Conceptually, a dependency-focused Mean-Time-To-Remediate (MTTR) metric can provide a historical perspective on how long it takes a given package to update vulnerable versions of its dependencies.
-However, existing MTTR metrics focus on a package fixing bugs in its own code, not its dependencies.
-Simultaneously, existing dependency update metrics do not aggregate values for the entire package and are not sensitive to aspects important for vulnerabilities (e.g., floating version constraints).
-The goal of this study is to aid industry practitioners, including developers, in assessing the risk of dependencies through a novel metric approximating mean-time-to-remediate vulnerabilities in their dependencies that is evaluated by an empirical study.
-We propose a novel algorithm for computing MTTR called MTTRdep and a companion metric called Mean-Time-To-Updatedep (MTTUdep), which considers all version updates, including vulnerability fix updates.
-We conduct a large-scale study using $163,207$ packages in npm, PyPI, and Cargo, of which only $22,513$ packages produce MTTRdep because of the lack of vulnerability data.
-We further study how package characteristics (e.g., contributors and version counts) influence MTTUdep and MTTRdep and explore how long packages retain outdated vulnerable dependencies in npm, PyPI, and Cargo.
-Our results indicate that industry practitioners can reliably use MTTUdep as a proxy for MTTRdep when available vulnerability data is insufficient.
+Industry practitioners are increasingly concerned with software that contains vulnerable versions of third-party dependencies that are included both directly and transitively.
+To address this problem, projects are encouraged to both (a) quickly update to non-vulnerable versions of dependencies and (b) be mindful of the update practices of the dependencies they choose to use.
+To this end, researchers have proposed metrics to measure the responsiveness of the development teams of the packages in keeping their dependencies updated: Mean-Time-To-Update (MTTU) and Mean-Time-To-Remediate (MTTR).
+While MTTU covers all dependencies, MTTR quantifies the time needed for a package to update its vulnerable dependencies.
+However, existing metrics fail to capture important nuances, such as considering floating versions and prioritizing recent updates, leading to inaccurate reflections of a development team’s update practices.
+The goal of this study is to aid practitioners in understanding how quickly packages update their dependencies.
+We propose two novel metrics, Mean-Time-To-Update for dependencies (MTTUdep) and Mean-Time-To-Remediate for vulnerable dependencies (MTTRdep), that overcome the limitations of existing metrics.
+We conduct an empirical study using $163,207$ packages in npm ($117,129$), PyPI ($42,777$), and Cargo ($3,301$) and characterize how the ecosystems differ in MTTUdep and MTTRdep, as well as what package characteristics influence MTTUdep and MTTRdep.
+We found that most packages have a relatively fast dependency update practice.
+We also found that older packages tend to have higher MTTUdep and MTTRdep values.
+We further study whether MTTUdep can be used as a proxy for MTTRdep when sufficient vulnerability data is not available.
+As we did not find enough statistical evidence for a strong proxy, our findings suggest that MTTUdep could only be partially used (may be used but with caution) as a proxy for MTTRdep when vulnerability data is not available.
+This latter finding is particularly important given that only 1363 npm (0.04%), 694 PyPI (0.11%), and 383 Cargo (0.20%) packages have reported vulnerabilities, and the existence of MTTUdep will allow practitioners to make more informed decisions about the dependencies they choose.
 1. Introduction
 Vulnerable dependencies are widely present in both open-source software (OSS) and proprietary codebases.
-According to the Synopsys 2025 “Open Source Security and Risk Analysis Report” (blackduck, ), 86% of codebases contain at least one vulnerable open source dependency, and 81% of codebases contain high or critical risk vulnerabilities resulting from dependencies.
+According to the Synopsys 2025 “Open Source Security and Risk Analysis Report” (blackduck, ), 86% of codebases contain at least one vulnerable open source dependency, and 81% of codebases contain high or critical risk externally reported vulnerabilities resulting from dependencies.
 The majority of codebases had vulnerable dependencies for more than two years despite the availability of a fixed version (ossf-scorecard, ; blackduck-2021, ).
 This delay occurs because of fear of breaking changes and the cost associated with updating vulnerable dependencies to fixed versions (derrKeepMeUpdated2017, ).
-Industry practitioners desire metrics that help assess the trustworthiness, risk, and security practices of open-source dependencies.
-For example, OpenSSF Scorecard (ossf-scorecard, ) considers 18 security practices to assess a package, including the “maintained” check, which determines if the project is maintained by checking activity in the last 90 days, and the “vulnerabilities” check detects if there are unfixed known vulnerabilities in the project or its dependencies.
-Mean-Time-To-Remediate (MTTR) is also often discussed in the context of vulnerability management (sonatype-2019, ).
-In contrast to Scorecard’s ‘maintained’ metrics, MTTR provides a historical perspective.
-However, MTTR has largely been discussed in the context of a project fixing its own vulnerabilities, not vulnerabilities in dependencies (gokhaleTimeStructureBased1999, ; calvoApplyingGoalQuestion2023, ; johnsonSurveySoftwareTools1988, ; morrisonMappingFieldSoftware2018, ).
-Dependency update metrics such as technical lag (gonzalez-barahona_technical_2017, ) consider management of dependency versions; however, they have not been applied to vulnerabilities.
-Conceptually, a lower value in a dependency update metric suggests faster and more consistent updates, signaling ongoing maintenance and reliability (sonatype-2024, ), which could help assess the security risks of adopting a package.
-However, the connection between dependency updates and exposure to vulnerabilities in dependencies has not yet been demonstrated.
-Practitioners assessing the risk associated with using a given dependency face two key challenges when considering how it manages vulnerabilities in its dependencies.
+This issue of vulnerable dependencies highlights the industry practitioners’ need for metrics to measure the responsiveness of development teams to updating their open-source dependencies.
+For example, the OpenSSF Scorecard (ossf-scorecard, ) evaluates a package based on 18 security practices.
+Among these practices are the “maintained” check, which determines if the project is maintained by checking activity in the last 90 days, and the “vulnerabilities” check, which detects if there are unfixed externally reported vulnerabilities in the project or its dependencies.
+While Scorecard’s “maintained” metric focuses on recent activity, other metrics, such as Mean-Time-To-Update(MTTU) and Mean-Time-To-Remediate (MTTR), provide a historical perspective on how long it takes for the development team of a package to update their dependencies.
+MTTU captures the time to update all dependency versions, while MTTR focuses specifically on the time to update vulnerable dependency versions to the fixed version (sonatype-2019, ).
+MTTR has been extensively discussed in the context of a project fixing its own vulnerabilities, but not vulnerabilities in dependencies (gokhaleTimeStructureBased1999, ; calvoApplyingGoalQuestion2023, ; johnsonSurveySoftwareTools1988, ; morrisonMappingFieldSoftware2018, ).
+Dependency update metrics, such as technical lag (gonzalez-barahona_technical_2017, ), consider management of dependency versions; however, they have not been applied to vulnerabilities.
+Furthermore, active maintenance (e.g., regular dependency update) is a desired criterion for developers in selecting a dependency as highlighted by Vargas et al.
+(larios_vargas_selecting_2020, ).
+In this context, a lower value in a dependency update metric suggests faster and more consistent updates, signaling ongoing maintenance and reliability (sonatype-2024, ).
+Practitioners measuring the responsiveness of a team in updating its vulnerable and outdated dependencies using dependency update metrics face two key challenges.
 1.
-Insufficient Vulnerability Data.
-Currently, only 2.4K of npm packages have known vulnerabilities, while the ecosystem hosts over 3.27M packages.
+Metric Limitations.
+The calculations used by existing dependency update metrics often do not handle floating version constraints or are not designed to localize vulnerabilities (coxMeasuringDependencyFreshness2015, ; gonzalez-barahona_technical_2017, ).
+Existing metrics also often calculate an update time for each of a dependency and for each version of a package, making it difficult to understand that package’s update practices as a whole.
+Additionally, existing metrics do not weight the recent dependency update practice which makes the existing metrics less actionable for developers.
+Finally, there are no existing publicly available dependency update metrics specifically for vulnerable dependencies.
+While some industry reports use MTTR in this way, their specific calculations are not available.
+2.
+Insufficient Externally Reported Vulnerability Data.
+A small fraction of packages in software ecosystems have reported vulnerabilities.
+For example, 1363 npm, 694 PyPI, and 383 Cargo (total 2.4K) packages have externally-reported vulnerabilities as of 2024-09-12.
 Practitioners can only compute MTTR for the dependents of those 2.4K packages.
-2.
-Metric Limitations.
-The calculations used by existing dependency update metrics often do not handle floating version constraints or are not designed to handle vulnerabilities (coxMeasuringDependencyFreshness2015, ; gonzalez-barahona_technical_2017, ).
-Existing metrics also often calculate an update time for each of a dependency and for each version of a package, making it difficult to understand that package’s update practices as a whole.
-Finally, there are no existing publicly available dependency update metrics specifically for vulnerabilities.
-While some industry reports use MTTR in this way, their specific calculations are not available.
-The goal of this study is to aid industry practitioners, including developers, in assessing the risk of dependencies through a novel metric approximating mean-time-to-remediate vulnerabilities in their dependencies that is evaluated by an empirical study.
-Specifically, we hypothesize that a package’s dependency update metric (Mean-Time-To-Update, MTTU) can serve as an indicator of how quickly the package would remediate security vulnerabilities (MTTR) in dependencies, even if the package never had a vulnerable dependency before.
-We also overcome challenges with dependency update calculations and propose novel algorithms for computing MTTU and MTTR, which we denote MTTUdep and MTTRdep, respectively.
-We use MTTUdep and MTTRdep to answer the following research questions: RQ1:Can MTTUdep be used as a proxy for MTTRdep?
-RQ2:How do package characteristics influence MTTUdep and MTTRdep?
-RQ3:How long do packages have outdated dependencies, and how long do packages have vulnerable dependencies?
-To answer these questions, we first collect package version release information and dependency relations of packages from the npm, PyPI, and Cargo ecosystems and compute MTTUdep and MTTRdep for $163,207$ packages.
-For RQ1, we perform an equivalence analysis using TOST (two one-side t-tests) and analyze correlations between MTTUdep and MTTRdep.
-For RQ1, we use correlation tests to understand the association between various package characteristics (e.g., contributors count, version count).
+0.49% npm, 0.83% PyPI, and 0.05% Cargo packages have directly depended on at least one of these 2.4k vulnerable packages in their lifetime.
+In addition, the lack of externally reported vulnerability data is also discussed by related research (alhazmi2007measuring, ; zahan_software_2023, ).
+The goal of this study is to aid practitioners in understanding how quickly packages update their dependencies through an empirical study using two novel dependency update metrics.
+With this goal, we conduct our study with four research questions.
+RQ1::
+How do we measure the MTTU and MTTR of a package including its dependencies such that the measure is responsive to more recent update practices and therefore actionable to the development team?
+We first propose novel algorithms for computing Mean-Time-To-Update (MTTU) and Mean-Time-To-Remediate (MTTR), which we denote MTTUdep and MTTRdep, respectively, to overcome challenges with existing dependency update metric calculations.
+After defining the two novel metrics, we performed an empirical analysis on npm, PyPI, and Cargo ecosystems.
+We collect package version release information and dependency relations of packages of $163,207$ packages from the three ecosystems and compute MTTUdep and MTTRdep.
+With the computed metrics, we answer the second research question.
+RQ2::
+How do packages in npm, PyPI, and Cargo differ in MTTU and MTTR?
+We analyze the distributions of MTTUdep and MTTRdep using violin plots to explore the differences in the ecosystems.
+In all three ecosystems, MTTRdep cannot be computed for 99.51% npm, 99.17% PyPI, and 99.95% Cargo packages since these packages have not depended upon any vulnerable direct dependencies in their lifetime.
+While dependency update metrics are useful, this lack of externally reported vulnerability data limits their application.
+To provide dependency update measurements for those 99%+ of packages without externally-reported vulnerabilities, we analyzed whether MTTUdep would provide a practical estimation/ proxy.
+This analysis is inspired by Dhrymes and Guerard (dhrymes1978introductory, ), who suggested that when a variable is unobservable, a proxy variable, which is a variable that can be used as a substitute for the missing one, can be used.
+This leads us to our third research question.
+RQ3::
+Can MTTU be used as a proxy for MTTR?
+We perform a proxy analysis, consisting of a set of statistical tests from literature, on MTTUdep and MTTRdep to explore if MTTUdep can serve as a proxy for MTTRdep.
+Finally, we would like to understand which package characteristics (e.g., contributors count, version count) have more influence on MTTUdep and MTTRdep, which leads to our last research questions.
+RQ4::
+How do package characteristics influence MTTU and MTTR?
+We use correlation tests to understand the association between nine package characteristics and MTTUdep/ MTTRdep.
 We further conduct regression analysis to quantify which package characteristics matter more in influencing MTTUdep and MTTRdep values.
-Finally, for RQ1, we analyze the distributions of MTTUdep and MTTRdep to explore how long packages retain outdated and vulnerable dependencies in the ecosystems.
 Contributions.
-In summary, this paper contributes (1) a detailed algorithm and process for quantifying the dependency update practice of a package using our novel metrics; (2) statistical hypothesis testing on using MTTUdep as a proxy for MTTRdep, when vulnerability data is not available; (3) a large-scale analysis of the dependency update metrics in npm, PyPI, and Cargo packages; (4) correlation and regression analysis illustrating which package characteristics impact the likelihood of higher MTTUdep and MTTRdep; and (5) actionable insights for open-source developers and supply chain security researchers based on our findings.
-We provide our replication package in Zenodo (zenodo-artifact, ) (available for reviewers till Oct’25). Upon acceptance of the paper, we will make it public.
-2. Related Work
-In this section, we provide a brief overview of the related work.
+In summary, this paper contributes (1) a detailed algorithm and process for quantifying the dependency update practice of a package using our novel metrics; (2) statistical hypothesis testing on using MTTU as a proxy for MTTR, when externally reported vulnerability data is not available; (3) a large-scale analysis of the dependency update metrics in npm, PyPI, and Cargo packages; and (4) correlation and regression analysis illustrating which package characteristics impact the likelihood of higher MTTU and MTTR.
+We provide our replication package in Zenodo (zenodo-artifact, ), currently restricted for reviewers only. Upon acceptance of the paper, we will make it public.
+2. Background And Related Work
+In this section, we provide a brief overview of the existing update metric and discuss related work.
 Technical Lag.
@@
 (zeroualiFormalFrameworkMeasuring2019, ) propose a formal framework for measuring technical lag in software ecosystems.
-They analyzed 4M releases of 500K npm packages considering the evolution of technical lag over time.
+They analyzed 4M releases of 500K npm packages, considering the evolution of technical lag over time.
 They found that technical lag induced by direct dependencies in npm packages increases over time due to missed updates, including major releases.
@@
 Although previous studies have explored technical lag in terms of general dependency updates, understanding vulnerable dependency update practice was not their goal.
-In contrast, we study the dependency update metric and vulnerable dependency update metric and investigate their relationship.
+In contrast, we study the dependency update metric and the vulnerable dependency update metric and investigate their relationship and other characteristics.
 Outdated and vulnerable dependencies.
@@
 They found that for most programming languages, a critical vulnerability persists on average for over a year before being fixed.
-Studies in outdated dependencies and vulnerable dependencies are focused on either all updates or only security updates but not both.
+Studies in outdated dependencies and vulnerable dependencies are focused on either all updates or only security updates, but not both.
 In contrast, we focus on both outdated and vulnerable dependencies and explore their relationship using our proposed metrics.
@@
 The metrics MTTU and MTTR have been used in the software reliability and maintenance domain for a long time (gokhaleTimeStructureBased1999, ; calvoApplyingGoalQuestion2023, ; johnsonSurveySoftwareTools1988, ; morrisonMappingFieldSoftware2018, ).
-Researchers have also studied different security metrics (time to close bug/vulnerability, window of exposure, vulnerability count) (davidsonGoodBadUgly2009, ; beresUsingSecurityMetrics2009, ) of various categories (time metric, vulnerability metric) in the software security domain, but these metrics are focused on packages having the vulnerability.
+Researchers have also studied different security metrics (time to close bug/vulnerability, window of exposure, vulnerability count) (davidsonGoodBadUgly2009, ; beresUsingSecurityMetrics2009, ) of various categories (time metric, vulnerability metric) in the software security domain, but these metrics are focused on vulnerabilities in packages [^1] Faults and bugs are considered in MTTR instead of vulnerability in Reliability domain research..
 In our study, we focus on measuring dependency update metrics for packages having vulnerable dependencies, not the vulnerable package itself.
@@
 $0.9.1:\operatorname{2022-03-17}$ $^{\wedge}2.48.0$ $2.48.0:\operatorname{2021-12-01}$ $3.3.0:\operatorname{2022-03-11}$ $100$
-$0.9.2:\operatorname{2022-03-22}$ ${}^{\wedge}3.4.0$ $3.4.0:\operatorname{2022-03-18}$ $3.4.0:\operatorname{2022-03-18}$ 00
+$0.9.2:\operatorname{2022-03-22}$ $^{\wedge}3.4.0$ $3.4.0:\operatorname{2022-03-18}$ $3.4.0:\operatorname{2022-03-18}$ 0
 3.1. Technical Lag Using tLag
-Time lag ($tLag$(zerouali_empirical_2018, ; decan_evolution_2018, ) is a way to measure technical lag to assess the outdatedness of a dependency in terms of time.
+Time lag (tLagtLag(zerouali_empirical_2018, ; decan_evolution_2018, ) is a way to measure technical lag to assess the outdatedness of a dependency in terms of time.
 Conceptually, $tLag$ measures the time difference between the release time of the latest version of a dependency and the release time of the version of the dependency used by a package, at the time the package was released.
-Let $pkg$ be a package, and $pkg_{v}$ indicate specific version $v$ of $pkg$.
-Let $dep$ be a direct dependency of $pkg$.
-When specifying $dep$ as a dependency, $pkg_{v}$ specifies a dependency constraint $pkg_{v}.dep.\bar{c}$.
+Let pkgpkg be a package, and $pkg_{v}$ indicate a specific version $v$ of pkgpkg.
+Let depdep be a direct dependency of pkgpkg.
+When specifying depdep as a dependency, $pkg_{v}$ specifies a dependency constraint $pkg_{v}.dep.\bar{c}$.
 Package managers use dependency constraints to resolve the highest version that satisfies the constraint.
-For example, a dependency constraint $pkg_{v}.dep.\bar{c}$ of “${}^{\wedge}1.2.3$” is satisfied for the range [¿=1.2.3, ¡ 2.0.0].
-We use the function $lastAllowed(dep,\bar{c},t)$ to denote the version resolution of $dep$ for constraint $\bar{c}$ at time $t$.
-We further use the function $latest(dep,t)$ to denote the latest version of $dep$ at time $t$ and $time(\cdot)$ to denote the time a given version is released.
-Hence, the $tLag$ of $pkg_{v}$ for $dep$ is formally defined as follows:
+For example, a dependency constraint $pkg_{v}.dep.\bar{c}$ of $^{\wedge}1.2.3$” is satisfied for the range [¿=1.2.3, ¡ 2.0.0].
+We use the function $lastAllowed(dep,\bar{c},t)$ to denote the version resolution of depdep for constraint $\bar{c}$ at time $t$.
+We further use the function $latest(dep,t)$ to denote the latest version of depdep at time $t$ and $time(\cdot)$ to denote the time a given version is released.
+Hence, the $tLag$ of $pkg_{v}$ for depdep is formally defined in Equation 1.
 (1) $tLag(pkg_{v},dep)=time(latest(dep,time(pkg_{v})))\\ -time(lastAllowed(dep,pkg_{v}.dep.\bar{c},time(pkg_{v})))$
@@
 The first two columns of Table 1 indicate the version with the date of different releases of package codemod-cli and the version constraint specified for dependency simple-git.
-The “lastAllowed(simple-git)” column, represents the resolved version of dependency simple-git with specified constraint at the time of the release of codemod-cli.
+The “lastAllowed(simple-git)” column represents the resolved version of dependency simple-git with specified constraint at the time of the release of codemod-cli.
 For computing the technical lag, we need to know the latest available version of the dependency simple-git at the time of the release of each version of codemod-cli according to Eq.
-1, which represents in column “latest(simple-git)”.
+1, which is represented in column “latest(simple-git)”.
 The tLag column is the computed technical lag.
@@
 Package-Level Metric.
-A dependency update metric should provide a way to aggregate all data points into one data for each package so that developers can make sense of the metrics and use them for comparison if desired.
-One of the findings of Wermke et al.
-(wermkeAlwaysContributeBack2023, ) was that developers’ dependency selection metrics and criteria focus on quickly accessible numbers and facts, such as downloads, GitHub stars, and time since last release, substantiating the necessity of having easily comprehensible metrics.
+A dependency update metric should combine all data points into one single value per package to make it easier for developers to understand and compare metrics.
 However, tLag computes the technical lag for each version of the package for a dependency as shown in Table 1.
 For developers, understanding what to do with all of these tLag values for a package with multiple releases and multiple dependencies is hard.
-Having an aggregation would improve the understandability of dependency update metrics.
+Having an aggregation would improve the utility of dependency update metrics.
 3
-Handling Vulnerabilities.
+Localizing Vulnerabilities.
 A security-oriented dependency update metric should incorporate the update characteristics of vulnerable dependencies.
+tLag includes bug fixes, feature updates, and security fixes as a whole.
 Technically, tLag’s measurement can be modified to only consider the dependency’s security fixes adopted by the package in its formal framework.
 However, no previous work measured tLag for measuring vulnerable dependency update practices.
-4. Proposed Dependency Update Metrics
-In this section, we first describe our design choices that fill the gaps presented in Section 3.2 for dependency update metrics measurement.
-Then, we provide an example of the nominal case using our metrics.
-We then provide a detailed methodology for defining the dependency update metrics in Section 4.3.
-4.1. High-Level Overview
+4
+Weighting Lifetime And Recent Practices.
+A dependency update metric should reflect more heavily the most recent dependency updates made by the package developers.
+However, tLag does not weight recency into its measurement.
+Recent update practice provides the most relevant information for ongoing maintenance of the package.
+For example, the development team’s update practice from 10 years ago should not be weighted as heavily as their current update practice.
+In addition, recency makes a dependency update metric more actionable since a metric with recency can be used to compare two packages with different lifespans (e.g., two packages with the same functionality, one with a shorter lifespan and the other with a longer lifespan).
+Actionability is a desired property for good software metrics (meneely_validating_metrics_2013, ).
+4. RQRQ1:: Novel Dependency Update Metrics
+In this section, we first describe our design choices that fill the gaps presented in Section 3.2 and lead to our novel metrics design (MTTUdep and MTTRdep) for RQRQ1:: .
+Then, we provide an example case study using our metrics.
+We then provide a detailed methodology for defining our novel dependency update metrics in Section 4.3.
+4.1. Intuition
 We design our metrics to only consider direct dependencies since the package only has control over which version to use for the direct dependencies.
-We begin by examining each of the $<$package, dependency$>$ relationships from a temporal perspective.
-We then split each $<$package, dependency$>$ relations into multiple intervals based on when the package or the dependency has a newer release (major, minor, or patch release).
+We begin by examining each of the <<package, dependency>> relationships from a temporal perspective.
+We then split each <<package, dependency>> relation into multiple intervals based on when the package or the dependency has a newer release (major, minor, or patch release).
 Our decision to split into intervals is to capture the benefits of floating version constraints without incentivizing this practice (Gap 1).
@@
 We do this to make sure that the dependency resolution accounts for the historical version releases of the dependency.
-After splitting into intervals, we mark “updated” as ‘false’ if the resolved dependency version for that interval does not match the highest available version of the dependency at that time and ‘true’ otherwise.
-Similarly, we mark “remediated” as ‘false’ if the resolved dependency version is vulnerable with a fixed version being available at the beginning of that interval and ‘true’ otherwise (Gap 3).
-We then aggregate the “updated” and “remediated” information of each $<$package, dependency$>$ to compute the MTTUdep and MTTRdep of each package (Gap 2).
-4.2. Nominal Case With Our Metrics
-Table 2. Running example of codemod-cli package with one of it’s depndency simple-git.
-row pkg pkg version dep dep constraint dep version dep highest rel. Interval start Interval end updated remediated
-1 codemod-cli $0.4.0$ simple-git ${}^{\wedge}1.130.0$ $1.132.0$ $2.13.1$ $\operatorname{2020-07-16}$ $\operatorname{2020-07-17}$ false true
-$\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$
-82 codemod-cli $0.8.6$ simple-git ${}^{\wedge}1.130.0$ $1.132.0$ $3.2.6$ $\operatorname{2022-03-03}$ $\operatorname{2022-03-05}$ false true
-83 codemod-cli $0.8.7$ simple-git ${}^{\wedge}2.48.0$ $2.48.0$ $3.2.6$ $\operatorname{2022-03-05}$ $\operatorname{2022-03-08}$ false true
-84 codemod-cli $0.9.0$ simple-git ${}^{\wedge}2.48.0$ $2.48.0$ $3.2.6$ $\operatorname{2022-03-08}$ $\operatorname{2022-03-11}$ false true
-85 codemod-cli $0.9.0$ simple-git ${}^{\wedge}2.48.0$ $2.48.0$ $3.3.0$ $\operatorname{2022-03-11}$ $\operatorname{2022-03-17}$ false false
-86 codemod-cli $0.9.1$ simple-git ${}^{\wedge}2.48.0$ $2.48.0$ $3.3.0$ $\operatorname{2022-03-17}$ $\operatorname{2022-03-18}$ false false
-87 codemod-cli $0.9.1$ simple-git ${}^{\wedge}2.48.0$ $2.48.0$ $3.4.0$ $\operatorname{2022-03-18}$ $\operatorname{2022-03-22}$ false false
-88 codemod-cli $0.9.2$ simple-git ${}^{\wedge}3.4.0$ $3.4.0$ $3.4.0$ $\operatorname{2022-03-22}$ $\operatorname{2022-03-29}$ true true
+After splitting into intervals, we mark as “updated”=false if the resolved dependency version for that interval does not match the highest available version of the dependency at that time and ‘true’ otherwise.
+Similarly, we mark as “remediated”=false if the resolved dependency version is vulnerable with a fixed version being available at the beginning of that interval and ‘true’ otherwise (Gap 3).
+We then aggregate the “updated” and “remediated” information of each <<package, dependency>> to compute the MTTUdep and MTTRdep of each package (Gap 2).
+This aggregation involves factoring in how old each interval is and weighting based on that.
+With our weighting mechanism, recent intervals get near full weight, and older intervals’ weight drops exponentially to zero (Gap 4).
+4.2. An Example With Our Metrics
+Table 2. Running example of codemod-cli package with one of its dependency simple-git.
+row
+pkg
+pkg
+version
+dep
+dep
+constraint
+dep
+version
+dep
+highest rel.
+Interval start Interval end
+Age Of Interval
+updated remediated
+1 codemod-cli $0.4.0$ simple-git $^{\wedge}1.130.0$ $1.132.0$ $2.13.1$ $\operatorname{2020-07-16}$ $\operatorname{2020-07-17}$ 1495 false true
 $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$
-In this section, we explain how to calculate the dependency update metrics. As a running example, we consider the dependency relation between codemod-cli and simple-git, as shown in Table 2.
+82 codemod-cli $0.8.6$ simple-git $^{\wedge}1.130.0$ $1.132.0$ $3.2.6$ $\operatorname{2022-03-03}$ $\operatorname{2022-03-05}$ 899 false true
+83 codemod-cli $0.8.7$ simple-git $^{\wedge}2.48.0$ $2.48.0$ $3.2.6$ $\operatorname{2022-03-05}$ $\operatorname{2022-03-08}$ 896 false true
+84 codemod-cli $0.9.0$ simple-git $^{\wedge}2.48.0$ $2.48.0$ $3.2.6$ $\operatorname{2022-03-08}$ $\operatorname{2022-03-11}$ 893 false true
+85 codemod-cli $0.9.0$ simple-git $^{\wedge}2.48.0$ $2.48.0$ $3.3.0$ $\operatorname{2022-03-11}$ $\operatorname{2022-03-17}$ 887 false false
+86 codemod-cli $0.9.1$ simple-git $^{\wedge}2.48.0$ $2.48.0$ $3.3.0$ $\operatorname{2022-03-17}$ $\operatorname{2022-03-18}$ 886 false false
+87 codemod-cli $0.9.1$ simple-git $^{\wedge}2.48.0$ $2.48.0$ $3.4.0$ $\operatorname{2022-03-18}$ $\operatorname{2022-03-22}$ 882 false false
+88 codemod-cli $0.9.2$ simple-git $^{\wedge}3.4.0$ $3.4.0$ $3.4.0$ $\operatorname{2022-03-22}$ $\operatorname{2022-03-29}$ 875 true true
+$\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$
+In this section, we explain how to calculate the dependency update metrics, Time-To-Update (TTUdep) and Time-To-Remediate (TTRdep), for a <<package, dependency>> relationship.
+As a running example, we consider the dependency relation between codemod-cli and simple-git, as shown in Table 2.
 We split the dependency relations into multiple intervals based on the release of a new version of the package “codemod-cli” or dependency “simple-git”.
@@
 In $V0.8.7$, codemod-cli updated the constraint for simple-git to $^{\wedge}2.48.0$.
-Even with this update, codemod-cli has not changed the constraint of simple-git to the available highest major version 3 and rather stayed at major version 2.
+Even with this update, codemod-cli has not changed the constraint of simple-git to the available highest major version 3, and rather stayed at major version 2.
 codemod-cli have $^{\wedge}2.48.0$ constraint for simple-git from $V0.8.7$ to $V0.9.1$.
-The false value in the “updated” column means that the package has an outdated version of the dependency in this interval.
+“updated”=false indicates that the package has an outdated version of the dependency in this interval.
 At the beginning of row 85, the resolved version of simple-git $V2.48.0$ was found vulnerable to four vulnerabilities (CVE-2022-24066 (CVE-2022-24066, ), CVE-2022-24433 (CVE-2022-24433, ), CVE-2022-25912 (CVE-2022-25912, ), and CVE-2022-25860 (CVE-2022-25860, )).
@@
 For this reason, we mark the intervals from rows 85-87 as “remediated” = false since codemod-cli has a vulnerable version of simple-git even though a fixed version is available.
-The false value in the “remediated” column means that the package has an outdated and vulnerable version of the dependency in this interval.
-In $V0.9.2$, codemod-cli bumped the constraint of simple-git to ${}^{\wedge}3.4.0$ which resulted in $V3.4.0$, a fixed version of the above vulnerability.
-So codemod-cli has the highest available version of simple-git in this interval (row 88) and so marked as true in the “remediated” column.
-When computing the Time-To-Update (TTUdep) for dependency relation between codemod-cli and simple-git, we sum up the intervals marked false in the “updated” column.
-For codemod-cli, the TTUdep is the delta between $\operatorname{2020-07-16}$ to $\operatorname{2022-03-22}$ (rows 1 - 87) which is $614$ days.
-Similarly, for computing Time-To-Remediate (TTRdep), we sum up the intervals marked false in the “remediated” column.
-For codemod-cli, the TTRdep is the delta between $\operatorname{2022-03-11}$ to $\operatorname{2022-03-22}$ (rows 85 - 87) which is $11$ days.
+“remediated”=false means that the package has an outdated and vulnerable version of the dependency in this interval.
+In $V0.9.2$, codemod-cli bumped the constraint of simple-git to $^{\wedge}3.4.0$, which resulted in $V3.4.0$, a fixed version of the above vulnerability.
+So codemod-cli has the highest available version of simple-git in this interval (row 88) and so marked “remediated”=true.
+When computing the Time-To-Update (TTUdep) for the dependency relation between codemod-cli and simple-git, we sum up the intervals with “updated”=false with exponential weighting.
+With the weighting factor, TTUdep for codemod-cli becomes 2.4 days, which is less than naively computing the delta between $\operatorname{2020-07-16}$ to $\operatorname{2022-03-22}$ (rows 1 - 87), which is $614$ days.
+Since this period of outdated dependency (rows 1 - 87) occurred in the year 2022, the weighting factor ensures giving less emphasis on that.
+Similarly, TTRdep for codemod-cli (rows with “remediated”=false) becomes 3.67 days.
+Because of the weighting factor, our computed TTRdep =3.67 is lower than naively summing up the delta between $\operatorname{2022-03-11}$ to $\operatorname{2022-03-22}$ (rows 85 - 87), which is $11$ days.
+Since this period of intervals with “remediated”=false is older, these intervals are weighted accordingly in TTRdep.
+We formally define MTTUdep and MTTRdep in Section 4.3 and explain our design choices.
 4.3. Metrics Definitions
 We present a formal definition of our proposed metrics in this section.
-Metric 1 ($\operatorname{\mathbf{Mean-Time-To-Update}}_{dep}$ : $\mathbf{MTTU}_{dep}$).
-MTTUdep of a package is the average aggregated time the package uses an outdated direct dependency version in its lifetime.
-We have described how to calculate the TTUdep for a $<$package, dependency$>$ relationship in Section 4.2.
-Formally speaking, TTUdep of a package $p_{i}$ with considering only dependency $p_{j}$ ($<p_{i},p_{j}>$ relationship) is:
-(2) ${TTU}_{dep}(p_{i},p_{j})=\sum\operatorname{intervals\ with\ updated=False}$
-MTTUdep of package $p_{i}$ with $n$ direct dependencies is:
+Metric 1 ($\operatorname{\mathbf{Mean-Time-To-Update}}_{dep}$ : $\mathbf{MTTU}_{dep}$)
+MTTUdep of a package is the weighted aggregated time the package uses an outdated direct dependency version in its lifetime.
+We have described how to calculate the TTUdep for a <<package, dependency>> relationship in Section 4.2.
+Formally speaking, TTUdep of a package $p_{i}$ with considering only dependency $p_{j}$ ($p_{i},p_{j}>$ relationship) is defined in Equation 2.
+(2) ${TTU}_{dep}(p_{i},p_{j})=\frac{\sum_{t}w_{t}d_{t}}{\sum_{t}w_{t}}$
+MTTUdep for package $p_{i}$ with $n$ direct dependencies is defined in Equation 3.
 (3) ${MTTU}_{dep}(p_{i})=\frac{\sum\limits_{j=1}^{n}{TTU}_{dep}(p_{i},p_{j})}{n}$
-Metric 2 ($\operatorname{\mathbf{Mean-Time-To-Remediate}}_{dep}$ : $\mathbf{MTTR}_{dep}$).
-MTTRdep of a package is the average aggregated time a package uses an outdated and vulnerable direct dependency version in its lifetime.
-We have described how to calculate the TTRdep for a $<$package, dependency$>$ relationship in Section 4.2.
-Formally speaking, TTRdep of a package $p_{i}$ with considering only dependency $p_{j}$ ($<p_{i},p_{j}>$ relationship) is:
-(4) ${TTR}_{dep}(p_{i},p_{j})=\sum\operatorname{intervals\ with\ remediated=False}$
-MTTRdep of package $p_{i}$ with $n$ direct historically vulnerable dependencies is:
+Here, $d_{t}$ indicates one interval duration with ‘updated’=false for $p_{i},p_{j}>$ relationship which ends at timestamp $t$.
+Also, $w_{t}=\exp(-\lambda a_{t})$ is the weight assigned for interval $d_{t}$ and $a_{t}$ is the age of the interval $d_{t}$.
+$\lambda$ is the decaying factor in this weighting function, and we set $\lambda=\frac{\ln(2)}{\tau}$.
+In this equation, $\tau$ is the half-life, and we set $\tau=2\ years$ since 2 years is used in literature to assess recent ongoing maintenance (miller_understanding_2025, ; li_comparison_2023, ).
+With this decaying weight, recent intervals approach full weight, while the weight of the older intervals decays exponentially to near zero.
+Our use of weighted average is inspired by similar other research (guzman_emotional_awareness_2013, ; ajiono_time_series_forecasting_2023, ; lee2016comparison, ; sikand_green_ai_2023, ).
+We considered linear ($w_{t}=max(a)-a_{t}+\epsilon$), exponential ($w_{t}=\exp(-\lambda a_{t})$), and inverse ($w_{t}=\frac{1}{a_{t}+\epsilon}$) as the choices for the weighting function based on the criteria described by Ulan et al.
+(ulan_weighted_2021, ) for weighted quality scoring for software metrics.
+We opted for the exponential weighting function since exponential weighting is more responsive to recent data than linear weighting, and is more configurable and robust than inverse weighting (using $\tau$).
+For example, if a development team wants to consider 3 years as an appropriate half-life for their specific case, they can configure the weighted version by changing $\tau=3\ years$.
+Metric 2 ($\operatorname{\mathbf{Mean-Time-To-Remediate}}_{dep}$ : $\mathbf{MTTR}_{dep}$)
+MTTRdep of a package is the weighted aggregated time a package uses an outdated and vulnerable direct dependency version in its lifetime.
+TTRdep of a package $p_{i}$ with considering only dependency $p_{j}$ is defined in Equation 4.
+(4) ${TTR}_{dep}(p_{i},p_{j})=\frac{\sum_{t}w_{t}d_{t}}{\sum_{t}w_{t}}$
+MTTRdep for $n$ direct dependencies is defined in Equation 5.
 (5) ${MTTR}_{dep}(p_{i})=\frac{\sum\limits_{j=1}^{n}{TTR}_{dep}(p_{i},p_{j})}{n}$
-5. Methodology
-In this section, we first present vulnerability, package metadata (versions and dependency relations), and package characteristics collection process.
-We then present a statistical testing process to verify if MTTUdep can be a proxy for MTTRdep (RQ1).
-We then present correlation tests and regression analysis to understand how package characteristics impact MTTUdep and MTTRdep (RQ1).
+Here, $d_{t}$ indicates one interval duration with ‘remediated’=false for $p_{i},p_{j}>$ relationship which ends at timestamp $t$.
+Also, $w_{t}=\exp(-\lambda a_{t})$ is the weight assigned for interval $d_{t}$ and $a_{t}$ is the age of the interval $d_{t}$.
+Takeaway 1: Our design of MTTUdep and MTTRdep overcomes the limitations of existing dependency update metrics.
+5. Empirical Study Methodology
+In this section, we first present vulnerability, package metadata (versions and dependency relations), and the packages’ characteristics collection process to apply the metric implementation in the three ecosystems (RQRQ2:: ).
+After that, we present a statistical testing process to verify if MTTU can be a proxy for MTTR (RQRQ3:: ).
+Lastly, we present correlation tests and regression analysis to understand how package characteristics impact MTTU and MTTR (RQRQ4:: ).
 5.1. Data Collection
-5.1.1. Vulnerability Information
+Vulnerability Information.
 We collect the CVE data for our chosen three ecosystems from osv.dev (osv-dev, ) for npm, PyPI, and Cargo packages on 2024-09-12.
 We rely on OSV since OSV aggregates CVE data from multiple sources (e.g., GitHub Security Advisories, PyPA, GoVulDB) (osv-data-sources, ) in one place.
-After downloading JSON formatted CVE data from OSV, we convert it into an SQL table that includes ecosystem, package name, CVE ID, version where the vulnerability was introduced, and version where the vulnerability is fixed.
-5.1.2. Package Metadata
+After downloading JSON-formatted CVE data from OSV, we convert it into an SQL table that includes ecosystem, package name, CVE ID, version where the vulnerability was introduced, and version where the vulnerability was fixed.
+Package Metadata.
 We collect the package-version data and dependency information for npm, PyPI, and Cargo packages from deps.dev on 2024-08-20, similar to other previous studies (lin_untrustide_2024, ; zahan_software_2023, ).
+We chose the three ecosystems to have a diverse set of ecosystems, where npm is the largest, PyPI is the oldest, and Cargo is the newest among the major software ecosystems.
 In our dataset, we have initially $2,603,314$ npm, $274,720$ PyPI, and $122,069$ Cargo packages.
-We collected data from deps.dev since it provides all the package versions and dependency information for our three chosen ecosystems.
-After data collection, we split each $<$package, dependency$>$ relation into multiple intervals.
-Since the dependency resolution can be complex and differs across ecosystems, we use deps.dev to perform the dependency resolution allowed by the dependency constraints with our added requirement of only using the available versions of the dependency before the interval start time.
-5.1.3. Package Characteristics
+We collected data from deps.dev since it provides all package versions and dependency information for our three chosen ecosystems.
+After data collection, we split each <<package, dependency>> relation into multiple intervals.
+Since the dependency resolution can be complex and differs across ecosystems, we use deps.dev to perform the dependency resolution allowed by the dependency constraints, with our added requirement (only using the available versions of the dependency before the interval start time).
+Package Characteristics.
 We examined several characteristics (from Saini et al.
 (saini_investigating_2020, )) for packages: the number of contributors; the number of dependencies and dependents; the number of version releases; the SourceRank score, and the number of forks and stars.
-On 2025-01-11, we downloaded these characteristics of each package from libraries.io to understand if these factors correlate with package’s MTTUdep and MTTRdep values.
+On 2025-01-11, we downloaded these characteristics of each package from libraries.io to understand if these characteristics influence a package’s MTTUdep and MTTRdep values.
 Libraries.io is used by other research as a data source (gu_self-admitted_2023, ; cao_towards_2023, ; he_migrationadvisor_2021, ).
 When counting the number of dependencies, we only use the number of dependencies of packages in their latest version.
-To ensure construct validity in downloading these factors, we manually inspected a sample of packages to verify the factors, and we found the downloaded factors to be accurate.
+To ensure construct validity in downloading these characteristics, we manually inspected a sample of packages to verify the characteristics, and we found the downloaded characteristics to be accurate.
 Additionally, we computed the number of major versions and package ages for our packages and added them to our analysis.
-5.2. Package Inclusion Exclusion Criteria
+5.2. Package Inclusion and Exclusion Criteria
 We begin with an initial dataset of $3,000,103$ ($2,603,314$ npm, $274,720$ PyPI, and $122,069$ Cargo) packages collected from deps.dev.
@@
 (miller_understanding_2025, ) used two years of residual activity followed by two years of no maintenance as criteria to find out the abandoned packages.
-Our criteria are inspired by Miller et al.
-since we want to include the packages that are maintained.
+Our criteria are inspired by Miller et al., since we want to include the packages that are maintained.
 Moreover, “two years” is a commonly used standard adopted by other research to measure whether a package is maintained or not (li_comparison_2023, ).
-However, our package selection criteria might miss packages that are less than two years old or have had no activity in the last two years (e.g., feature complete packages (coelho2017modern, )) even if they are not abandoned.
-We then used our exclusion criteria: a package without any dependency should be excluded.
+However, our package selection criteria might miss packages that are less than two years old or have had no activity in the last two years (e.g., feature complete packages (coelho2017modern, )), even if they are not abandoned.
+We then used our exclusion criteria: a package without any dependencies should be excluded.
 Since our metrics characterize packages’ dependency update practice, packages without any dependencies do not fit into our study.
-The number of packages after this exclusion criteria is $163,207$ ($117,129$ npm, $42,777$ PyPI, and $3,301$ Cargo packages).
+The number of packages after these exclusion criteria is $163,207$ ($117,129$ npm, $42,777$ PyPI, and $3,301$ Cargo packages), which will be the final set we use in our analyses.
 Out of these $163,207$ packages, $22,513$ ($17,263$ npm, $5,158$ PyPI, and $92$ Cargo) packages have at least one vulnerable dependency in their lifetime, supporting our initial motivation of lack of vulnerability data.
@@
 We also create indexes on the most frequently accessed keys to speed up data retrieval for our metrics calculations.
-In the database, we compute if the resolved dependency version matches the highest available version of the dependency during each interval and mark each interval accordingly, as shown in the “updated” column in Table.
-2.
-Similarly, we then compute if the resolved dependency version for each interval is vulnerable to any security advisory even though a fixed version is available and mark the interval accordingly in the “remediated” column in Table.
-2.
-5.4. Statistical and Correlation Tests
-To answer RQ1, we claim that if MTTUdep and MTTRdep are statistically equivalent, then MTTUdep can serve as a proxy for MTTRdep.
-Since we did not expect our samples to follow a normal distribution, we initially selected the non-parametric Mann-Whitney test.
-However, we also applied the t-test to ensure a more conservative analysis, as our data is not ordinal.
-Both tests were conducted using the Python library scipy, and the Wilcoxon test was also implemented using the same package.
-To assess normality, we used the Kolmogorov-Smirnov test.
-Our initial hypotheses focused on the equivalence of data distributions.
-For these, following the approach of Papotti et al.
-(papotti_acceptance_2024, ), we employed the TOST (Two One-Sided Test) approach, originally proposed by Schuirmann et al.
-(schuirmann_hypothesis-testing_1981, ) and commonly used in pharmacological and food sciences to evaluate equivalence within a specified range (guidance_statistical_2001, ; meyners_equivalence_2012, ).
-Other research also used the TOST test in their analysis (shelton_rosita_2021, ; labunets2018no, ; labunets2017graphical, ; palheiros2024does, ).
+In the database, we compute if the resolved dependency version matches the highest available version of the dependency during each interval and mark each interval accordingly, as shown in the “updated” column in Table 2.
+Similarly, we then compute if the resolved dependency version for each interval is vulnerable to any security advisory, even though a fixed version is available, and mark the interval accordingly in the “remediated” column in Table 2.
+5.4. Proxy Analysis
+To answer RQRQ3:: , we apply some statistical tests on MTTUdep and MTTRdep data to analyze whether a variable (MTTUdep) can be used as a substitute/ proxy for another (MTTRdep) from the literature.
+(1) TOST (Two One-Sided Test).
+Schuirmann et al.
+(schuirmann_hypothesis-testing_1981, ) proposed an equivalence test, known as TOST (Two One-Sided Test), in bioequivalence studies to determine if a treatment (e.g., a drug) can be used as a substitute for another.
+This method then was used in pharmacological/ food science (guidance_statistical_2001, ; meyners_equivalence_2012, ), medical research (mascha2011equivalence, ), and later also adopted to software engineering and security (papotti_acceptance_2024, ; shelton_rosita_2021, ; labunets2018no, ; labunets2017graphical, ; palheiros2024does, ).
 According to TOST, two distributions, $x$ and $y$, are considered equivalent if $x\cdot\delta<y<x\cdot 1/\delta$, where $\delta=0.8$.
-The TOST method can use either Mann-Whitney test or t-test as its underlying direction test.
-We reported the results of the t-test because it was more conservative than the Mann-Whitney test.
-Additionally, to measure the association between MTTUdep and MTTRdep, we calculated both the Pearson correlation coefficient and Spearman’s rank correlation coefficient (similar to Cox et al.
+We report TOST with Mann-Whitney tests as the underlying difference tests.
+(2) Regression with Wald Test.
+Several works (wooldridge2010econometric, ; mahnken2014evaluating, ) proposed the use of regression to identify proxy.
+A proxy should have (1) a statistically significant slope ($p_{value}$ ¡ 0.05); (2) a normal distribution for random error with a mean of zero and small variance.
+Additionally, Montgomery et al.
+(montgomery2000measuring, ) used (3) a Wald test on the proxy variable coefficient.
+Based on these, we ran an Ordinary Least Squares (OLS) regression using the statsmodel library in Python, with a Wald test.
+(3) Sensitivity Analysis.
+Seltzer (seltzer2021perilous, ) proposed sensitivity analysis for proxy analysis, with $GFI$ (Goodness-of-Fit Index), $AGFI$ (Adjusted Goodness-of-Fit Index), and $NFI$ (Normal Fit Index).
+The indices should be ¿ 0.90 for an acceptable fit (marsh1996assessing, ).
+We ran the sensitivity analysis using the semopy library in Python.
+(4) Correlation Analysis.
+High correlation is one indication of a proxy (frost1979proxy, ).
+We calculated both the Pearson correlation coefficient (as used by Oh et al.
+(oh2019identification, )) and Spearman’s rank correlation coefficient (similar to Cox et al.
 (cox_measuring_2015, )).
-High ($>0.7$) or moderate coefficient values ($>0.5$) would suggest a strong to moderate positive relationship (ratner2009correlation, ).
-Such findings would also support the notion that MTTUdep could effectively serve as a proxy for MTTRdep.
+High (>0.7>0.7) or moderate coefficient values (>0.5>0.5) would suggest a strong to moderate positive relationship (ratner2009correlation, ).
+We utilize the scipy.stats package from Python to calculate these correlations.
 5.5. Package Characteristics Analysis
-For analyzing eight package-level characteristics in RQ1, we use correlation tests to estimate if there is any correlation between these characteristics and packages’ corresponding MTTUdep and MTTRdep values.
+For analyzing nine package-level characteristics in RQRQ4:: , we use correlation tests to estimate if there is any correlation between these characteristics and packages’ corresponding MTTUdep and MTTRdep values.
 Specifically, we test a hypothesis about the association between the following independent variables and their continuous dependent variables (MTTUdep and MTTRdep):
-Contributors Count ($H_{1}$). We hypothesize packages with fewer contributors are less likely to have updated dependencies since they have less capacity for maintenance and dependency management.
-Dependents Count ($H_{2}$).
-We hypothesize packages with fewer dependents are less likely to have updated dependencies.
+($H_{1}$) Contributors Count..
+We hypothesize that packages with fewer contributors are less likely to have updated dependencies since they have less capacity for maintenance and dependency management.
+($H_{2}$) Dependents Count.
+We hypothesize that packages with fewer dependents are less likely to have updated dependencies.
 Fewer dependents may lead to fewer feature updates, fewer bug fixes, and fewer version releases, which in turn may result in less updated dependencies.
-Dependency Count ($H_{3}$).
-We hypothesize packages with fewer dependencies are more likely to have updated dependencies.
+($H_{3}$) Dependency Count.
+We hypothesize that packages with fewer dependencies are more likely to have updated dependencies.
 We expect fewer dependencies to be more manageable, and thus, these projects may have more updated dependencies.
-Version Count ($H_{4}$).
-We hypothesize packages with fewer number of version releases are less likely to have updated dependencies.
-Since these packages have less available versions, they may pay less attention to their dependency management.
-Major Version Count ($H_{5}$).
-We hypothesize packages with fewer major version releases are more likely to have updated dependencies.
+($H_{4}$) Version Count.
+We hypothesize that packages with fewer version releases are less likely to have updated dependencies.
+Since these packages have fewer available versions, they may pay less attention to their dependency management.
+($H_{5}$) Major Version Count.
+We hypothesize that packages with fewer major version releases are more likely to have updated dependencies.
 Since these packages have fewer major versions to maintain, they may pay more attention to their dependency management.
-Package Age ($H_{6}$).
-We hypothesize packages with lower ages are more likely to have updated dependencies.
+($H_{6}$) Package Age.
+We hypothesize that packages with lower ages are more likely to have updated dependencies.
 Lower package age may mean the development team is more proactive in dependency management since the package is not old or mature enough.
 Lower package age may also mean that the dependencies do not have scope for publishing many newer versions, which in turn may mean more updated dependencies in these packages.
-SourceRank ($H_{7}$).
+($H_{7}$) SourceRank.
 The Package SourceRank score indicates the package quality, popularity, and community metrics calculated in libraries.io dataset (saini_investigating_2020, ; sun_using_2023, ).
 This metric depends on several factors, such as the presence of a README file, license, following SEMVER, recent updates, and the number of contributors.
-We hypothesize packages with lower SourceRank scores are less likely to have updated dependencies.
+We hypothesize that packages with lower SourceRank scores are less likely to have updated dependencies.
 Since these packages are of low quality, they may not have a lot of dependents or contributors, which in turn may result in less updated dependencies.
-Forks And Stars Count ($H_{8}$).
-We hypothesize packages with fewer forks and stars are less likely to have updated dependencies.
-Fewer forks and stars may mean fewer people are using and looking into these packages, which in turn may mean less activity, less version releases, and less updated dependencies in these packages.
+($H_{8}$) Forks Count.
+We hypothesize that packages with fewer forks are less likely to have updated dependencies.
+Fewer forks may mean fewer people are using and looking into these packages, which in turn may mean less activity, fewer version releases, and fewer updated dependencies in these packages.
+($H_{9}$) Stars Count.
+We hypothesize that packages with fewer stars are less likely to have updated dependencies.
+Fewer stars may indicate fewer people are using and looking into these packages, which consequently may suggest less activity, fewer version releases, and fewer updated dependencies.
 5.6. Regression Analysis
-Correlation results indicate whether two independent variables are associated with each other or not.
-However, correlation does not establish causation.
-To identify such a cause-and-effect relationship using mathematical modeling, we opt for applying a technique of multilinear regression analysis.
-The multilinear regression model takes independent variables (e.g., package characteristics) and one dependent variable (MTTUdep or MTTRdep) and gives results on the relationship between each independent variable and the dependent variable.
-This analysis provides valuable insights on predicting MTTUdep or MTTRdep using the package’s characteristics.
-This analysis also results in $p$-values and coefficients, indicating which package characteristics might have a significant impact on MTTUdep or MTTRdep.
-This process will produce a $p$-value for each independent variable to indicate whether the relationship between this variable and the outcome is statistically significant.
-Finally, we consider the variables with $p$-value $<0.05$ as significant variables, and the package characteristics associated with these variables are identified as key package characteristics.
-We use statsmodels package of Python to conduct the correlation tests and to build the regression model.
-6. Results
-6.1. RQ1: Can MTTUdep serve as a proxy for MTTRdep?
-We statistically assess the relationship between MTTUdep and MTTRdep to determine if MTTUdep can effectively serve as a proxy for MTTRdep in  RQ1.
-We used TOST (Two One-Sided Test) to investigate the equivalence of MTTUdep and MTTRdep.
-The single one-sided tests are done with the Mann-Whitney test.
+Correlation analysis measures pairwise relationships between two variables, but does not account for potential interactions between multiple variables.
+In contrast, a multilinear regression model allows evaluating the combined effect of all independent variables in predicting the dependent variable while controlling for the others.
+Moreover, strong relationships between two variables from correlation analysis might be influenced by the presence of other variables (confounding variables).
+A multilinear regression model controls for confounding effects, providing an understanding of the unique contribution of each independent variable in predicting the dependent variable.
+So, for RQRQ4:: , we use multilinear regression models that take independent variables (e.g., package characteristics) and one dependent variable (MTTUdep or MTTRdep) and give results on the relationship between each independent variable and the dependent variable.
+This analysis provides valuable insights into predicting MTTUdep or MTTRdep using the package’s characteristics.
+This analysis also results in $p_{value}$s and coefficients, indicating which package characteristics might have a significant impact on MTTUdep or MTTRdep.
+This process will produce a $p_{value}$ for each independent variable to indicate whether the relationship between this variable and the outcome is statistically significant.
+To control for family-wise type-I error inflation due to testing multiple dependent models (e.g., models that share the same dependent variable) together, we applied the Bonferroni correction (bonferroni1935, ) for the $p_{value}$ threshold to determine the statistical significance level.
+Since we test nine different characteristics, according to Bonferroni correction, the $p_{value}$ of a test needs to be <0.05/9or 0.0055<0.05/9\ or\ 0.0055 to be considered as significant ($p<0.0055$).
+Package characteristics with a significant test result are identified as key package characteristics.
+We use the statsmodels package of Python to conduct the correlation tests and to build the regression model.
+6. Empirical Study Results
+6.1. RQRQ2:: : How do packages in npm, PyPI, and Cargo differ in MTTU and MTTR?
+In RQRQ2:: , we empirically analyze the MTTUdep and MTTRdep metrics for the three ecosystems.
+We choose a violin plot instead of box plot since a violin plot shows everything a box plot shows, e.g., medians, ranges, variability, and the violin plot’s shape shows the density of the data similar to a density estimation plot (hintzeViolinPlotsBox1998, ).
+Fig.
+3 shows a violin plot of MTTUdep in npm, PyPI, and Cargo.
+All the plots are right-skewed, indicating that most packages have a low MTTUdep.
+For instance, $50\%$ npm package has MTTUdep of less than $51$ days.
+Also, every plot has a long tail, indicating that every ecosystem has some packages that do not update their dependencies for a long time (max MTTUdep =2653=2653 days).
+In addition, interquartile ranges are small and comparable for the three ecosystems (Cargo with $1\sim 39$ days, npm with $4\sim 45$ days, and PyPI with $11\sim 54$ days).
+Fig.
+3 shows a violin plot visualizing the distribution across MTTRdep in days for Cargo, npm, and PyPI packages.
+The overall pattern is similar to MTTUdep.
+The relative distribution width and the right-skewed nature of MTTRdep are similar to MTTUdep.
+Interquartile ranges in MTTRdep are comparable for the three ecosystems (Cargo with $6\sim 24$ days, npm with $10\sim 42$ days, and PyPI with $12\sim 45$ days), similar to MTTUdep.
+A smaller interquartile range indicates MTTUdep and MTTRdep data is less spread and less variable.
+We could compute MTTUdep for $163,207$ ($117,129$ npm, $42,777$ PyPI, and $3,301$ Cargo) packages, and MTTRdep for $22,513$ ($17,263$ npm, $5,158$ PyPI, and $92$ Cargo) packages.
+This corroborates our initial motivation for conducting this study: the lack of vulnerability data.
+Figure 2. ${MTTU}_{dep}$
+Figure 3. ${MTTR}_{dep}$
+Comparison to Prior Work.
+In contrast to prior research (kula_developers_2018, ), we found that the majority of packages in an ecosystem have a lower MTTRdep.
+The reason behind our metrics encompassing lower MTTRdep values is that our metrics weight recency, and older dependency update practice has less impact on our metrics.
+We also observed a long-tail distribution for both MTTUdep and MTTRdep in all three ecosystems.
+A significant number of packages lag behind in keeping dependencies updated, which is similar to the observations of Cox et al.
+(cox_measuring_2015, ).
+In short, regularly updating dependencies might not be a widespread practice.
+Our observation is similar to vulnerable dependency updates as well.
+Even with weighting, some packages took months or longer to remediate known vulnerable dependencies, similar to observations from Kula et al.
+(kula_developers_2018, ).
+This suggests that challenges in keeping dependencies updated are not specific to any ecosystem but rather general in nature.
+Takeaway 2: Most packages in npm, PyPI, and Cargo have relatively fast dependency update practices. The small interquartile ranges indicate consistent dependency update practice with each ecosystem.
+Table 3. TOSTs Results
+Ecosystem MTTUdep $\times 0.8$ << MTTRdep << MTTUdep $\div 0.8$
+All $U=$2.5\text{.}{10}^{08}$\$, $p=0.02$ $U=$1.9\text{.}{10}^{08}$\$, $p<0.01$
+Cargo $U=4668$, $p=0.51$ $U=2806$, $p=$7.8\text{.}{10}^{-5}$\$
+npm $U=$1.5\text{.}{10}^{08}$\$, $p=0.04$ $U=$1.2\text{.}{10}^{08}$\$, $p=$6.1\text{.}{10}^{-278}$\$
+PyPI $U=$1.3\text{.}{10}^{07}$\$, $p=0.0005$ $U=$1.0\text{.}{10}^{07}$\$, $p=$5.6\text{.}{10}^{-103}$\$
+Table 4. Proxy Analysis Result
+Criteria Result Pass?
+TOST Statistically significant Yes
+Regression - Coefficient Statistically significant Yes
+Regression - $R^{2}$ Moderate Moderately
+Regression - Wald test Statistically significant Yess
+Sensitivity - GFI <0.90<0.90 No
+Sensitivity - AGFI <0.90<0.90 No
+Sensitivity - NFI <0.90<0.90 No
+Correlation - Pearson Moderate and positive Moderately
+Correlation - Spearman Moderate and positive Moderately
+6.2. RQRQ3:: : Can MTTU serve as a proxy for MTTR?
+As we specified in Section 5.4, the proxy analysis covered four tests/ analyses:
+(1) TOST (Two One-Sided Test).
 The results of the TOSTs are reported in Table 4.
-We also visualized the TOST for MTTUdep and MTTRdep from all three ecosystems in Fig.
-2.
-Table 3. TOSTs Results
-Ecosystem MTTUdep ×$\times 0.8$ $<$ MTTRdep $<$ MTTUdep ÷$\div 0.8$
-[…]
-[…]
-[…]
-[…]
-The green area corresponds to $x\cdot\delta<y<x\cdot 1/\delta$ where $\delta=0.8$, $x$ is MTTUdep, and $y$ is MTTRdep.
-Figure 2. Distribution behind TOST MTTUdep vs. MTTRdep.
-The results of the TOSTs show statistical equivalence between MTTUdep and MTTRdep ($p<0.05$) within the given range.
-Further, we also ran correlation analyses between MTTUdep and MTTRdep, in which both Spearman and Pearson correlation coefficients are $0.5$.
-The correlation analysis supports the results of our TOST tests as we can conclude that MTTUdep and MTTRdep are moderately and positively correlated since all correlation coefficients are $>=0.5$.
-This result confirms that as the MTTUdep increases for a package, it is highly probable that the MTTRdep also increases (and vice versa), supporting the use of MTTUdep as a proxy for MTTRdep when we do not have enough vulnerability data to compute MTTRdep.
-Takeaway 1: MTTUdep can be used as a proxy for MTTRdep since MTTUdep and MTTRdep are statistically equivalent (and moderately and positively correlated) in all three observed ecosystems.
-6.2. RQ1: How do other package characteristics influence MTTUdep and MTTRdep?
-Figure 3. Correlation matrix on packages’ characteristics and MTTUdep and MTTRdep values.
-6.2.1. Correlation Analysis.
-We tested eight package-level characteristics that might influence MTTUdep and MTTRdep.
-We present the correlation matrix with the packages’ characteristics and MTTUdep and MTTRdep data in Figure 3.
+The results of the TOSTs show statistical equivalence between MTTUdep and MTTRdep ($p<0.05$), except for Cargo ($p>0.05$).
+(2) Regression with Wald Test.
+The coefficient of MTTUdep from the OLS regression returned positive and also statistically significant ($\beta=0.69,SE=0.056,t=12.21,p<0.001$), which indicates that, on average, when MTTUdep increases by one unit, MTTRdep also increases by 0.69.
+The $R^{2}$, however, shows that MTTUdep only explained 30.5% of the MTTRdep’s variance ($R^{2}=0.305$).
+The interpretation of this $R^{2}$ varies in different domains.
+If we refer to Hair et al.
+(hair2022partial, ), we can interpret this $R^{2}$ as a moderate explanatory power.
+The Wald test also returns statistically significant ($F=149.1,p<0.001$), which confirms that MTTUdep contributes to explaining MTTRdep.
+(3) Sensitivity Analysis.
+The sensitivity analysis returns $GFI=0.83,AGFI=0.67$, and $NFI=0.83$.
+As these indices are lower than the acceptable threshold (0.90), this result indicates that MTTUdep does not provide a good fit to MTTRdep.
+(4) Correlation Analysis.
+Pearson correlation coefficients returns 0.552 with $p=$1.12\text{.}{10}^{-28}$\$.
+Spearman, on the other hand, returns 0.689 with $p=$1.47\text{.}{10}^{-49}$\$.
+The correlations show that MTTUdep and MTTRdep are moderately and positively correlated.
+This correlation shows that MTTUdep scales linearly with MTTRdep.
+The results from the four tests/analyses are summarized in Table 4.
+In summary, (1) MTTUdep is statistically equivalent to MTTRdep; (2) MTTUdep increases along with the increases of MTTRdep, with a moderate explanatory power; (3) MTTUdep does not provide an acceptable fit for MTTRdep; and (4) MTTUdep is moderately and positively correlated with MTTRdep.
+Considering all metrics as distinct criteria, MTTUdep strongly satisfied three criteria, moderately satisfied three others, and failed to meet the other three left, suggesting partial adequacy as a proxy for MTTRdep.
+This suggested partial adequacy indicates that developers may use MTTUdep as a proxy for MTTRdep with caution and encourages further research to find better proxies for MTTRdep.
+Takeaway 3: MTTUdep can only partially serve as a proxy for MTTRdep. This suggests that MTTUdep may be used as a proxy for MTTRdep with caution, and future research is needed to find a better proxy.
+6.3. RQRQ4:: : How do package characteristics influence MTTU and MTTR?
+Correlation Analysis.
+We present the correlation matrix with the nine packages’ characteristics and MTTUdep and MTTRdep data in Figure 4.
 The correlation heatmap provides insight into the relationships between various factors and their potential impact on MTTUdep and MTTRdep.
@@
 Contributors Count ($H_{1}$): We hypothesized that packages with fewer contributors would be less likely to maintain updated dependencies.
-From the heatmap, we can see that the contributors count show minimal correlations with MTTUdep (-0.04) and MTTRdep (-0.07), suggesting a limited influence on dependency management.
+From the heatmap, we can see that the contributors count shows minimal correlations with MTTUdep (-0.07) and MTTRdep (-0.07), suggesting a limited influence on dependency management.
 Dependents Count ($H_{2}$): We hypothesized that packages with fewer dependents are less likely to have updated dependencies.
-The results indicate a negligible correlation between dependents count and MTTUdep or MTTRdep.
-This suggests that the number of dependents does not play a substantial role in determining the timeliness of dependency updates.
-Dependency Count ($H_{3}$): We hypothesized packages with fewer dependencies to have more updated dependencies.
-The results show almost no correlation between dependency count and MTTUdep or MTTRdep.
-This implies that the number of dependencies in a package is not associated with dependency management efficiency.
+The results indicate a weak negative correlation between dependents count with MTTUdep (-0.17) and MTTRdep (-0.16).
+This suggests that the number of dependents plays a limited role in determining the timeliness of dependency updates.
+Dependency Count ($H_{3}$): We hypothesized that packages with fewer dependencies would have more updated dependencies.
+The results show a weak negative correlation between dependency count with MTTUdep (-0.10) and MTTRdep (-0.11).
+This implies that the number of dependencies in a package has little association with dependency management efficiency.
 Version Count ($H_{4}$): We hypothesized that packages with fewer version releases are less likely to have updated dependencies.
-However, the correlation matrix revealed no meaningful relationship between the number of version releases and either MTTUdep or MTTRdep.
-This suggests that the frequency of version releases does not significantly impact how quickly dependencies are updated.
+However, the correlation matrix revealed a moderate negative relationship between the number of version releases and MTTUdep (-0.55) and MTTRdep (-0.41).
+This supports our hypothesis that packages with more version releases are likely to have updated dependencies.
+Figure 4. Correlation matrix on packages’ characteristics and MTTUdep and MTTRdep values.
 Major Version Count ($H_{5}$): We hypothesized that packages with fewer major version releases are more likely to have updated dependencies.
-However, the correlation matrix revealed no meaningful relationship between the number of major version releases and either MTTUdep or MTTRdep.
-This suggests that the number of available major versions to maintain does not significantly impact how quickly dependencies are updated.
+However, the correlation matrix revealed a weak negative relationship between the number of major version releases and MTTUdep (-0.16) and MTTRdep (-0.14).
+This suggests that the number of available major versions to maintain has a limited impact on how quickly dependencies are updated.
 Package Age ($H_{6}$): We hypothesized that packages with lower ages are more likely to have updated dependencies.
-However, the correlation matrix invalidated this hypothesis with a weak correlation score ($0.32$).
-This suggests that as packages become mature with age, they tend to keep dependencies less updated.
-However, this conclusion should be treated with caution since our metrics might be age-sensitive (more discussion on this in Section 7).
+However, the correlation matrix reveals MTTUdep (0.10) and MTTRdep (0.07) have a weak correlation with package age.
+This suggests that package age has little impact on having updated dependencies.
+This also strengthens our intuition of creating weighted versions, MTTUdep and MTTRdep, to eliminate the potential age-sensitivity of MTTUdep and MTTRdep.
 SourceRank ($H_{7})$.
 We hypothesized that low quality packages (e.g., packages with lower SourceRank scores) are less likely to have updated dependencies.
-The correlation matrix supports this hypothesis for MTTRdep with a weak correlation score ($-0.22$), which means low-quality packages have higher MTTRdep values.
-However, low quality packages do not necessarily have higher MTTUdep values ($-0.06$ coefficient).
-Forks And Stars Count ($H_{8}$): We hypothesized that packages with fewer forks and stars are less likely to have updated dependencies.
-The results reveal weak to moderate positive correlations between stars (0.65) and forks (0.72) with maintainers count, indicating that more popular packages may have more active development.
-However, their relationships with MTTUdep and MTTRdep remain insignificant, suggesting that forks and stars do not directly translate into faster dependency updates.
-Overall, our results indicate weak or negligible correlations between the tested factors and dependency management efficiency as measured by MTTUdep and MTTRdep.
+The correlation matrix shows limited support for this hypothesis since low-quality packages do not necessarily have weak correlation with MTTUdep (0.09-0.09) and MTTRdep (0.10-0.10).
+Forks Count ($H_{8}$): We hypothesized that packages with fewer forks are less likely to have updated dependencies.
+The results reveal a weak to moderate positive correlation between forks (0.86) with maintainers count, indicating that more forked packages may have more active development.
+However, the relationships between forks count with MTTUdep and MTTRdep remain insignificant, suggesting that having more forks does not directly translate into faster dependency updates.
+Stars Count ($H_{9}$): We hypothesized that packages with fewer stars are less likely to have updated dependencies.
+There is a weak to moderate positive correlation between stars and maintainers count (0.85), indicating that more popular packages may have more active development.
+However, the relationships between star count with MTTUdep and MTTRdep are insignificant, implying that more stars do not directly indicate faster dependency updates.
+Overall, our results indicate weak or negligible correlations between the tested factors and MTTUdep and MTTRdep.
 These findings suggest that the analyzed factors may not strongly influence how updated dependencies are, and further exploration of other variables or non-linear relationships is recommended.
-Takeaway 2: Package characteristics (except package age and SourceRank score) have negligible correlations with dependency updatedness.
-Older packages are associated with higher MTTUdep, and popular and better quality packages (higher SourceRank score) are associated with lower MTTRdep.
-6.2.2. Regression Analysis (MTTUdep As Dependent Variable).
-Interpreting the multilinear regression model’s characteristics, we found 0.073 as R-squared value when MTTUdep was the dependent variable.
-R-squared value 0.073 indicates that 7.3% variation in the dependent variable can be explained by the model.
-Although the model is statistically significant (p-value ¡ 0.05 from the F-test), lower R-squared value indicates that other factors, beyond the nine dependent variables, also substantially impact MTTUdep.
-The F-statistic of this model is large (1399), and Prob(F-statistic) is 0.00.
+Takeaway 4: Package characteristics (except version count) have negligible correlations with dependency updatedness. Packages with higher version count are associated with lower MTTUdep and MTTRdep.
+Regression Analysis (MTTUdep As Dependent Variable).
+Interpreting the multilinear regression model’s characteristics, we found 0.039 as the $R^{2}$ value when MTTUdep was the dependent variable.
+$R^{2}$ value 0.039 indicates that 3.9% variation in the dependent variable can be explained by the model.
+Although the model is statistically significant ($p_{value}$ ¡ 0.0055 from the F-test), a lower $R^{2}$ value indicates that other factors, beyond the nine dependent variables, also substantially impact MTTUdep.
+The F-statistic of this model is large (547.7), and Prob(F-statistic) is 0.00.
 This indicates that at least one of the independent variables (or predictors) has a non-zero relationship with MTTUdep.
-We then look into the coefficients and p-values associated with each of the independent variables.
+We then look into the coefficients and $p_{value}$s associated with each of the independent variables.
 The coefficient indicates the expected change in the dependent variable for a one-unit change in one independent variable while holding other independent variables constant.
-The p-value associated with the t-statistic indicates whether the independent variable is statistically significant in explaining the variation in the dependent variable.
-A lower p-value (¡ 0.05) would indicate statistical significance.
-We have found that some independent variables have positive coefficients, which indicates an increase in one of these variables would result in an increase in our dependent variable, MTTUdep.
-Independent variables with positive coefficients are contributors count (0.0204), stars (0.0035), dependencies count (2.3027), and package age (0.1257).
-The dependencies count has the largest positive coefficient, which means that packages with more dependencies need longer times to update their dependencies.
-In addition, older packages tend to have longer MTTUdep as indicated by the 0.1257 coefficient, justifying our hypothesis ($H_{3}$) on dependencies count.
-We also found some independent variables have negative coefficients.
+The $p_{value}$ associated with the t-statistic indicates whether the independent variable is statistically significant in explaining the variation in the dependent variable.
+A lower $p_{value}$ (¡ 0.0055) would indicate statistical significance.
+We have found that some independent variables have positive coefficients, which indicates that an increase in one of these variables would result in an increase in our dependent variable, MTTUdep.
+Independent variables with positive coefficients are contributors count (0.0025), dependents count ($5.3\text{.}{10}^{05}$), forks (0.0002), and package age (0.0114).
+The package age has the largest positive coefficient, which indicates older packages tend to have longer MTTUdep, justifying our hypothesis ($H_{6}$) on package age.
+However, the rest of the coefficients are small, which indicates these independent variables do not have a significant impact on MTTUdep.
+We also found that some independent variables have negative coefficients.
 Negative coefficients indicate that an increase in that independent variable would result in a decrease in the dependent variable, MTTUdep.
-Independent variables with negative coefficients are dependents count (-0.0013), forks (-0.0085), SourceRank (-24.3203), and major versions count (-0.9068).
-SourceRank, which indicates package quality and popularity, shows strongly negative coefficients.
+Independent variables with negative coefficients are SourceRank (-2.2260), stars (-0.0002), dependencies count (-0.0493), and major versions count (-0.4766).
+SourceRank, which indicates package quality and popularity, shows moderate negative coefficients.
 Higher SourceRank indicates an overall better project.
-The model predicts if SourceRank for a package increases by 1 (moving to a “better” rank), MTTUdep drops by 24 days.
+The model predicts that if SourceRank for a package increases by 1 (moving to a “better” rank), MTTUdep drops by 2 days.
 Moreover, packages that have released more major versions tend to have a lower MTTUdep.
-Finally, all dependent variables except dependent count were found to be statistically significant (p-value ¡ 0.05).
-Takeaway 3: Packages with a lot of dependencies or older packages are more likely to have higher MTTUdep.
-Also, popular and better quality packages (with higher SourceRank scores) are more likely to have lower MTTUdep.
-6.2.3. Regression Analysis (MTTRdep As Dependent Variable).
-For MTTRdep as the dependent variable, the multilinear regression model results in 0.097 as R-squared value.
-R-squared = 0.097 indicates that the model explains 9.7% of the variation in MTTRdep.
-A lower R-squared value indicates that other unmodeled factors play an important role in determining MTTRdep, which is similar to our observations in MTTUdep.
-The F-statistic of this model is 246.8, and Prob(F-statistic) is 0.00.
+Finally, independent variables, except dependent count and forks, were found to be statistically significant ($p_{value}$ ¡ 0.0055).
+Takeaway 5: Older packages are more likely to have higher MTTUdep. Also, popular and better quality packages (with higher SourceRank scores) are more likely to have lower MTTUdep.
+Regression Analysis (MTTRdep As Dependent Variable).
+For MTTRdep as the dependent variable, the multilinear regression model results in 0.026 as the $R^{2}$ value.
+$R^{2}$ = 0.026 indicates that the model explains 2.6% of the variation in MTTRdep.
+A lower $R^{2}$ value indicates that other unmodeled factors play an important role in determining MTTRdep, which is similar to our observations in MTTUdep.
+The F-statistic of this model is 65.4, and Prob(F-statistic) is $2.5\text{.}{10}^{-119}$.
 This indicates the model is statistically significant and at least one of the dependent variables (or predictors) has a non-zero relationship with MTTRdep.
-We then look into the coefficients and p-values associated with each of the independent variables.
-We found that forks, stars, SourceRank, dependencies count, major version count and package age are statistically significant (p-value ¡ 0.05).
-On the other hand, the contributors count, and dependents count are not statistically significant (p-value ¿ 0.05) in modeling MTTRdep.
-As a predictor, forks (coefficient 0.0070) and stars (coefficient 0.0014) indicate a positive but small effect.
-More forks and stars are associated with slightly higher MTTRdep.
-Dependencies count (coefficient 0.5028), and package age (coefficient 0.1239) show a moderate positive effect on MTTRdep.
-Additionally, the major version count (coefficient -2.4954) shows a moderate negative effect, whereas SourceRank (coefficient -34.8183) shows a strong negative effect on MTTRdep.
+We then look into the coefficients and $p_{value}$s associated with each of the independent variables.
+We found that independent variables, except for the dependent count, forks, and stars, are statistically significant ($p_{value}$ ¡ 0.0055) in modeling MTTRdep.
+As a predictor, forks (coefficient 0.0008), dependents count (coefficient $2.3\text{.}{10}^{-05}$ and package age (coefficient 0.0091) indicate a positive but small effect.
+Older packages are associated with slightly higher MTTRdep.
+Additionally, the major version count (coefficient -0.3453), dependencies count (coefficient -0.1213), contributors count (coefficient -0.0107), and SourceRank (coefficient -1.3284) show a moderate negative effect on MTTRdep.
 This observation is similar to our previous observation with MTTUdep.
-According to the model’s prediction, an increase in SourceRank by 1 (i.e., a higher SourceRank indicates a better package) would result in a 34.8-day reduction in MTTRdep.
-Takeaway 4: Similar to MTTUdep, older packages are more likely to have higher MTTRdep.
+According to the model’s prediction, an increase in SourceRank by 1 (i.e., a higher SourceRank indicates a better package) would result in a 1-day reduction in MTTRdep.
+Takeaway 6: Similar to MTTUdep, older packages are more likely to have higher MTTRdep.
 Also, popular and better quality packages (with higher SourceRank scores) are more likely to have lower MTTRdep.
 Packages with higher major versions are more likely to have lower MTTRdep.
-6.3. RQ1: How long do packages have outdated dependencies, and how long do packages have vulnerable dependencies?
-In RQ1, we empirically analyze the MTTUdep and MTTRdep metrics for the three ecosystems.
-We choose violin plot instead of box-plot since violin plot shows everything box-plot shows, e.g., medians, ranges, variability, and violin plot’s shape shows the density of the data similar to a density estimation plot (hintzeViolinPlotsBox1998, ).
-Figure 4. ${MTTU}_{dep}$
-Figure 5. ${MTTR}_{dep}$
-Fig.
-4 shows a violin plot of MTTUdep in npm, PyPI, and Cargo.
-All the plots are right-skewed, indicating that most packages have a low MTTUdep.
-For instance, $50\%$ npm package has MTTUdep of less than $174$ days.
-In terms of distribution width, Cargo has a compact distribution, and npm is the most widely spread (Cargo ¡ PyPI ¡ npm).
-Also, every plot has a long tail, indicating that every ecosystem has some packages that do not update their dependencies for a long time (max MTTUdep $=4296$ days).
-Fig.
-5 shows a violin plot visualizing the distribution across MTTRdep in days for Cargo, npm, and PyPI packages.
-The first, second (mean), and third quantiles are shown as , , and , respectively.
-The overall pattern is similar to MTTUdep.
-The relative distribution width and the right-skewed nature of MTTRdep is similar to MTTUdep.
-Cargo and PyPI have concentrated distributions, with most MTTRdep (e.g., interquartile range) being less than $219$ and $281$ days.
-Conversely, npm shows a more spread distribution than Cargo and PyPI, indicating variability in MTTRdep in npm.
-We could compute MTTUdep for $163,207$ ($117,129$ npm, $42,777$ PyPI, and $3,301$ Cargo) packages, and MTTRdep for $22,513$ ($17,263$ npm, $5,158$ PyPI, and $92$ Cargo) packages.
-This corroborates our initial motivation for conducting this study, the lack of vulnerability data.
-Takeaway 5: Even without potentially abandoned packages, a subset of packages show extreme delays in keeping their dependencies and vulnerable dependencies updated.
-Comparison to Prior Work.
-Similar to prior research (kula_developers_2018, ), we also found that the majority of packages in an ecosystem do not keep dependencies updated.
-We also observed a long-tail distribution for both MTTUdep and MTTRdep in all three ecosystems.
-A significant number of packages lag behind in keeping dependencies updated, which is similar to the observations of Cox et al.
-(cox_measuring_2015, ).
-In short, regularly updating dependencies is not a widespread practice.
-Our observation is similar to vulnerable dependency updates as well.
-Many packages took months or longer to remediate known vulnerable dependencies, similar to observations from Kula et al.
-(kula_developers_2018, ).
-This suggests that challenges in keeping dependencies updated are not specific to any ecosystem but rather general in nature.
 7. Discussions And Implications
+7.1. Practical Implications For Developers
 When MTTUdep And MTTRdep Should Be Used.
-Our study provides new insights by establishing the relationship between MTTUdep and MTTRdep.
-We found that although a package’s MTTUdep is moderately correlated with its MTTRdep, MTTUdep and MTTRdep are statistically equivalence.
+Our study provides new insights by exploring the relationship between MTTUdep and MTTRdep.
+While MTTUdep does not fully meet all the criteria to serve as a proxy for MTTRdep, it satisfies six out of nine criteria, which can be considered as a partial proxy.
+This finding provides enough evidence for MTTU to be a practical indication of MTTR, especially when externally reported vulnerability data is not available.
 In practical terms, packages that are slow to update dependencies also tend to be slow in updating vulnerable dependencies.
-This finding implies that improving general dependency update practice will likely improve vulnerable dependency update practice, which seems logical, but we provide statistical evidence supporting our hypothesis.
-Having said that, developers should use MTTRdep when available and only MTTUdep as a proxy for MTTRdep when MTTRdep is not available.
-Package Age’s Influence In MTTUdep and MTTRdep.
-Package age is weakly correlated with MTTUdep and has a negligible correlation with MTTRdep in our analysis.
-In our regression analysis, we have used package age as an independent variable with MTTUdep and MTTRdep as dependent variables.
-However, since we design our metrics based on a temporal perspective, our metrics could have an inherent age sensitivity.
-So, conclusions like ‘older packages tend to have slightly higher MTTUdep’ should be treated cautiously.
-Since older packages might have more data points (negligible but positive correlation with version count, major version count, and dependencies count), package age could influence our analysis.
-Future research should explore how to model the package age in MTTUdep and MTTRdep (e.g., one way could be to normalize the metrics by package age or by only taking the last $n$ years into account).
-Influence of Unmodeled Package Characteristics.
-Prior work shows that developers hesitate to update dependencies due to a fear of breaking changes, a lack of awareness or knowledge about available updates, and sometimes a lack of motivation to invest time in updates (derrKeepMeUpdated2017, ).
-These human factors are the likely reasons behind the long tail of MTTUdep and MTTRdep data.
-Our findings also support this.
-Packages with many contributors or higher dependents count do not have a better MTTUdep or MTTRdep, in contrast to our hypothesis.
-A package used by thousands of other packages is not guaranteed to have dependencies updated.
-Both multilinear regression models in our study present an R-squared value of less than 10%.
-A lower R-squared value indicates that other unmodeled factors substantially affect MTTUdep and MTTRdep.
-The development team dynamic, the use of pinning and floating, the cost and efforts needed in testing, and the size of the codebase could be such possible unmodeled factors.
-Practical implications for developers.
-We confirmed that improving general dependency update frequency likely improves security as well.
+This finding implies that improving general dependency update practice will likely improve vulnerable dependency update practice as well.
+Having said that, developers should use MTTRdep when available and only MTTUdep as a proxy for MTTRdep with caution when MTTRdep is not available.
+Use Floating-Minor With Regular Major Updates.
+Our results indicate that improving general dependency update frequency likely improves security as well.
 This should be a strong incentive for the developers to treat dependency updates as an important part of routine maintenance, not an afterthought.
 Even scheduling periodic dependency update sprints or using automated update notifications to stay on top of the new releases can be an effective strategy.
-Allowing the latest version of the dependency using the dependency specification (e.g., * or $latest$) is the best way to keep MTTUdep and MTTRdep low (even zero).
+Allowing the latest version of the dependency using the dependency specification (e.g., * or latestlatest) is the best way to keep MTTUdep and MTTRdep low (even zero).
 However, that might not be possible due to the issues with breaking changes (ochoa_breaking_2022, ; raemaekers_semantic_2014, ; venturini_i_2023, ).
@@
 Our recommendation relies on developers using SEMVER correctly (pinckney_large_2023, ; li_large-scale_2023, ; dietrich_dependency_2019, ) so that the benefits of floating versions can be leveraged.
-Gaming Metrics.
+Dependency Selection Criteria.
+Our findings suggest that packages with higher SourceRank scores showed somewhat more efficient update practice.
+In practice, this means a popular, actively maintained package is more likely to receive timely updates (and even external contributions) than an obscure one.
+When choosing a dependency, a well-maintained package should be prioritized if that serves the required functionality.
+In addition, our findings suggest that packages with a higher number of versions and a higher number of major versions are likely to have better practices for updating vulnerable dependencies.
+A higher number of versions and major versions indicates the development team of the package has released additional features, is active in maintenance, and is more responsive to making changes.
+In case of a vulnerability in dependencies, this development team might be more prompt in releasing a new version of the package with mitigating the vulnerability, or they use a version constraint (e.g., floating) which automatically adopts security fixes from their dependencies.
+7.2. Practical Implication For Researchers
+Prior work shows that developers hesitate to update dependencies due to a fear of breaking changes, a lack of awareness or knowledge about available updates, and sometimes a lack of motivation to invest time in updates (derrKeepMeUpdated2017, ).
+These human factors are the likely reasons behind the long tail of MTTUdep and MTTRdep data.
+Our findings also support prior research, revealing that packages with many contributors or higher dependents count do not have a better MTTUdep or MTTRdep, in contrast to our hypothesis.
+Even packages with thousands of dependents are not guaranteed timely updates for their own dependencies.
+Researchers should focus more on exploring the human factors to uncover ways to balance the cost and benefit of dependency updates.
+In addition, both multilinear regression models in our study present an $R^{2}$ value of less than 10%.
+A lower $R^{2}$ value indicates that other unmodeled factors substantially affect MTTUdep and MTTRdep.
+The development team dynamics, the use of pinning and floating, the cost and efforts needed in testing, and the size of the codebase could be such possible unmodeled factors.
+Researchers should explore further if such other unmodeled factors influence MTTUdep and MTTRdep.
+7.3. Practical Implication For Tool Builders
+In this study, we provide an extensive evaluation with our proposed novel dependency update metrics.
+Tool builders can incorporate our dependency update metrics into dependency management tools for developers to make them more accessible.
+Wermke et al.
+(wermkeAlwaysContributeBack2023, ) found that developers’ dependency selection metrics and criteria focus on quickly accessible numbers and facts, such as downloads, GitHub stars, and time since last release, substantiating the necessity of having easily comprehensible metrics.
+To this end, our metrics give quickly accessible a single number for a package that is quickly accessible and actionable for developers.
+Similarly, security risk assessment tools (e.g., OpenSSF Scorecard) could also incorporate our dependency update metrics in their assessment to help developers better assess the security risk of a package.
+7.4. Gaming Metrics
 While it might seem that using a loose floating constraint and bumping dependency constraint periodically could “game” our metrics, the underlying outcome is that a package spends minimal time with outdated dependency versions and ensures that security fixes are adopted automatically.
-By doing so, this strategy reduces the risk of exploitation from using an outdated dependency version with known security vulnerability, such as log4Shell (everson_log4shell_2022, ; rahman_less_2024, ).
+By doing so, this strategy reduces the risk of exploitation from using an outdated dependency version with a known security vulnerability, such as log4Shell (everson_log4shell_2022, ; rahman_less_2024, ).
 In short, trying to game our metrics by adopting floating version constraints is essentially reducing this attack vector for a package.
@@
 This strategy might make the packages susceptible to malicious package updates (xz-incident (williams_research_2025, )).
-Since floating version constraints allow automatic updates, it can also facilitate propagating malicious package updates.
-Influence of Dependencies Count in MTTUdep and MTTRdep.
-In our design, we compute MTTUdep and MTTRdep by dividing the summation of TTUdep and TTRdep by the number of dependencies and number of historically vulnerable dependencies, respectively, in Eq.
-3 and  5.
-A package with a high number of dependencies has more room for having a better value in MTTUdep while keeping the majority of dependencies updated.
-However, MTTRdep is not influenced by dependencies count since we only consider the amount of time spent with vulnerable dependency versions and the number of historically vulnerable dependencies.
+Since floating version constraints allow aucan also facilitate propagating malicious package updates.
 8. Threats to Validity
-In this section, we discuss potential limitations and how they might impact the interpretation of our results.
 External Validity.
 The main external validity threat is the generalizability of our results to characterize other ecosystems.
-We carefully chose the three ecosystems to have a diverse set of ecosystems, where npm is the largest, PyPI is the oldest, and Cargo is the newest among the major software ecosystems.
 While each ecosystem possesses unique features that might not directly correlate with those we studied, we believe the insights gained should also be broadly applicable to other ecosystems.
@@
 Additionally, we do not consider whether the vulnerable dependency version is exploitable or reachable (cisa-vex, ) from the package.
-We treat all vulnerabilities equally regardless of the CVSS score or the severity of the vulnerability.
+We treat all vulnerabilities equally, regardless of the CVSS score or the severity of the vulnerability.
+Using the severity of vulnerabilities as a weighting factor in our metrics would be an interesting future work.
 After downloading the data from deps.dev, we manually checked 20 packages’ versions and relations with the public package registries and found that the data is accurate.
@@
 Our analysis omits package versions not adhering to SEMVER rules, a conservative choice to enable a more rigorous analysis.
-In addition, Open Source Insights dependency resolution fails in some cases (e.g.
-missing timestamp), and we mark those as warnings in our dataset.
+In addition, Open Source Insights dependency resolution fails in some cases (e.g., missing timestamp), and we mark those as warnings in our dataset.
 We do not calculate update metrics for those packages, and we argue that this might have a very small impact on our results.
@@
 In this study, we introduced two dependency update metrics, MTTUdep and MTTRdep, to quantify the updatedness of dependencies in open-source software packages.
-Our large-scale empirical analysis across the npm, PyPI, and Cargo ecosystems demonstrated that MTTUdep can serve as a proxy for MTTRdep when vulnerability information is unavailable.
+Our large-scale empirical analysis across the npm, PyPI, and Cargo ecosystems demonstrated that MTTUdep can serve partially as a proxy for MTTRdep when vulnerability information is unavailable.
 Furthermore, our statistical analysis highlighted the relationships between package characteristics and dependency update behavior, providing actionable insights for developers, maintainers, and software supply chain researchers.
@@
 Expanding this analysis to other ecosystems, combining with transitive dependencies, and incorporating qualitative insights from developers could further refine our understanding of dependency updatedness.
+Data Availability
+The code and data for the analysis in this paper are all available in our replication package in Zenodo (zenodo-artifact, ).
+It is currently restricted for reviewers only, but we will make it public upon acceptance.